### v1

# DefendCLI: {Command-Line} Driven Attack Provenance Examination

Abstract
Endpoint Detection and Response (EDR) solutions embrace the method of attack provenance graph to discover unknown threats through system event correlation. However, this method still faces some unsolved problems in the fields of interoperability, reliability, flexibility, and practicability to deliver actionable results. Our research highlights the limitations of current solutions in detecting obfuscation, correlating attacks, identifying low-frequency events, and ensuring robust context awareness in relation to command-line activities. To address these challenges, we introduce DefendCLI, an innovative system leveraging provenance graphs that, for the first time, delves into command-line-level detection. By offering finer detection granularity, it addresses a gap in modern EDR systems that has been overlooked in previous research. Our solution improves the precision of the information representation by evaluating differentiation across three levels: unusual system process calls, suspicious command-line executions, and infrequent external network connections. This multi-level approach enables EDR systems to be more reliable in complex and dynamic environments. Our evaluation demonstrates that DefendCLI improves precision by approximately $1.6\times$ compared to the state-of-the-art methods on the DARPA Engagement Series attack datasets. Extensive real-time industrial testing across various attack scenarios further validates its practical effectiveness. The results indicate that DefendCLI not only detects previously unknown attack instances, which are missed by other modern commercial solutions, but also achieves a $2.3\times$ improvement in precision over the state-of-the-art research work.
Index Terms:
Endpoint Detection and Response, Attack Investigation, Advanced Persistent Threat.
I Introduction
Traditional security measures, like anti-virus software and firewalls, frequently fall short in detecting contemporary threats that employ advanced evasion strategies, such as memory injection and fileless attacks. These limitations highlight the need for advanced Endpoint Detection and Response (EDR) solutions to detect complex and stealthy attacks, providing better insights into behavior and operations to improve security incident detection. An essential improvement of modern EDR systems is the attack provenance graph, which visually represents the relationships between different system events to discover potential advanced threats.
While there is growing interest in provenance-based EDR systems, their adoption in real-world scenarios remains limited, and comprehensive evaluation is still lacking. To address this gap, we deployed a provenance-based EDR [1] to automatically analyze system audit logs from the Microsoft Cyber Defense Operations Center (CDOC) and passed the results to security analysts for verification. During this process, analysts encountered challenges in efficiently utilizing the provenance-based results and we summarized four key unresolved research questions (RQs) that require further investigation.
Figure 1: Motivation Feedback: A provenance-based detector [1] deployed in an enterprise environment. Only 0.07% of the original alerts were truly actionable, providing accurate attack-related malicious activities, such as command-line executions and network connections, for a valid verification.
1
RQ1-Interoperability: How can we reduce the complexity of attack provenance graphs to provide more actionable results? Provenance graphs are often overly complex and cluttered with invalid information, making interpretation difficult and misaligned with analysts’ priorities.
Malicious activities often blend seamlessly into legitimate system operations, causing provenance-based EDRs to generate an overwhelming number of anomalies that lack actionable insights, as illustrated in Fig 1 and Fig 2. While these results may not necessarily be false alarms, they fail to provide definitive evidence of an intrusion, making manual investigation even more challenging and hindering efforts to trace the root cause and focus on suspicious activities.
These challenges highlight the need for a more refined approach to provenance graph construction, which can minimize complexity, enhance detection accuracy, and deliver more relevant, actionable insights for security analysts.
Figure 2: Omission: Even if a process chain is malicious, but it cannot be analyzed if we do not known what command-line are executed, of which effective threat verification is affected.
2
RQ2-Reliability: How can we provide a more effective approach for anomaly detection on different operating system platforms? Existing anomaly detection methods [2, 3, 4] often adopt a one-size-fits-all approach, failing to account for OS-specific characteristics. While this generalization increases versatility, it overlooks critical real-world complexities.
Enterprise security policies often segment networks into distinct zones, such as the demilitarized zone (DMZ) and internal office networks. The DMZ primarily hosts externally facing services, like DNS and web servers that typically run Linux, whereas internal office networks support end-user devices, predominantly Windows-based. Each segment presents unique behavioral patterns, attack surfaces, and vulnerabilities, making OS-specific anomaly detection essential for effective threat identification.
A more tailored approach, one that considers the distinct operational norms and security risks of each OS, can significantly improve the accuracy and efficiency of EDR solutions. By refining detection mechanisms to align with platform-specific behaviors, EDR systems can reduce false positives, enhance threat visibility, and provide more actionable results.
3
RQ3-Flexibility: How can we tailor AI-based detection mechanisms to minimize false positives and improve adaptability in dynamic environments? The adoption of AI in security detection and response is accelerating due to its superior capability in identifying zero-day threats compared to traditional signature-based systems.
However, AI-driven solutions introduce significant challenges, particularly an elevated false positive rate in dynamic environments where factors such as new device deployments, software updates, and network reconfigurations frequently trigger erroneous alerts.
Moreover, the inherent “black-box” nature of AI complicates security investigations, as opaque decision-making processes undermine interpretability, trust, and operational transparency.
To address these limitations, AI models should incorporate adaptive learning mechanisms that accommodate environmental variability while maintaining a stringent false positive control. Enhancing explainability through interpretable AI techniques and leveraging contextual awareness in anomaly detection can further improve precision, thereby fostering greater trust and operational efficacy in security operations.
4
RQ4-Practicability: How can we evolve beyond binary decision-making frameworks in EDR systems to enable more granular and nuanced threat analysis? Traditional security frameworks that classify alerts as strictly malicious or benign are increasingly inadequate, as modern cyber threats often blur the lines between legitimate and malicious activity. Advanced attack techniques exploit these gray areas, rendering binary classifications ineffective in identifying sophisticated threats.
To address this limitation, AI-driven security solutions are shifting toward threat-scoring models that assess incidents based on severity, providing a more nuanced analysis. However, many of these systems remain rule-based, necessitating frequent updates to adapt to evolving attack patterns. This continuous maintenance imposes a significant operational burden on security vendors, limiting scalability and responsiveness to emerging threats.
To enhance efficiency, a more automated, adaptive and actionable security orchestration framework is required - one that minimizes manual intervention while dynamically refining threat assessments. By integrating machine learning-driven anomaly detection, contextual analysis, and automated incident response, EDR systems can improve detection precision, reduce operational overhead, and offer a more intelligent, scalable approach to threat management.
Our Contributions. To address the aforementioned RQs, DefendCLI presents a practical solution that, for the first time, delves into the command-line-level detection within the attack provenance graph, improving accuracy while avoiding unrelated information.
Command-line interactions encompass both interactive shell operations and OS API executions, providing direct visibility into system-level behaviors. Despite attackers’ attempts at command-line obfuscation, sophisticated multi-stage attacks require numerous system function executions, inevitably producing traceable command-line artifacts that cannot be fully concealed from detection systems. Despite the widespread adoption of command-line monitoring in rule-based EDR systems, its formal representation in attack provenance graphs and systematic integration with AI-driven anomaly detection frameworks remains underdeveloped.
Our solution employs an explainable AI system to match the known malicious command-line with well-defined rule sets and detect unknown command-line activities with the SimHash-Based [5] anomaly detection mechanism. Furthermore, the overall solution focuses on multi-level attack provenance examination, including unusual system process calls, suspicious command-line executions, and infrequent external network connections, which can resist defense evasion to some extent.
To provide a more automated security orchestration scheme, we integrate Large Language Models (LLMs) into our threat recommendation system to generate rapid preliminary assessments, enabling efficient triage before comprehensive threat analysis. Extensive experiments and evaluations show that DEFENDCLI effectively manages real-world cyber threats by identifying previously undetected attack instances. Our contributions are summarized as follows:
  • Interoperability: We introduce a novel attack provenance graph structure specifically tailored for command-line-driven, multi-level differentiation detection. This refined structure reduces the complexity associated with traditional provenance graphs, resulting in more precise information representation and improved actionable insights for security analysts.
  • Reliability: Our approach incorporates an explainable anomaly detection framework, integrating defense evasion signatures and known attack patterns from both Linux and Windows platforms into an AI-driven architecture. The system also employs graph algorithms for multi-level provenance analysis alongside SimHash-based differentiation of unknown command-line executions, significantly reducing false positives by moving away from pre-training unreliable normal behavior profiles.
  • Flexibility: The framework supports dynamic adaptation through configurable hyperparameters, optimizing performance for various operational environments. By leveraging a retrieval-augmented generation (RAG) approach with LLMs, we generate prioritized incident assessments with detailed explanations, named snapshot alarms, thus streamlining alert triage and mitigating analyst alert fatigue.
  • Practicability: We evaluate the performance of our system against state-of-the-art research work and modern commercial EDR tools on academic benchmark datasets and industrial real-time detection. The result shows the practical applicability and superior effectiveness of our solution for modern offensive and defensive confrontation.
II Related Work
The attack provenance graph is a powerful and comprehensive tool that maps the origins, evolution, and interconnections of entities and events within a system, with a particular emphasis on data flow and processes [6, 7]. By constructing detailed network diagrams, it traces the journey of data, illustrating how information is accessed, transmitted, and modified across various endpoints.
We argue that, despite significant progress made by existing methods in leveraging provenance analysis for system monitoring, their effectiveness in detecting and explaining command-line activities remains limited. These activities, however, often provide actionable insights for threat confirmation, which are closely related to the RQs discussed above.
This section provides a comparative analysis of related work by discussing their limitations across four key dimensions [8] that significantly impact the effectiveness of the command-line activity detection: restricted effectiveness obfuscation detection (OD), unclear and insufficient attack correlation (AC), incapable of low-frequency events detection (LFD), and limited context awareness (CA).
Table I qualitatively compares DefendCLI with previous methods in their granularity levels to detect command-line activities, based on their respective algorithmic designs (AD). A more rigorous quantitative evaluation is provided in Section V for the evidential validation.
TABLE I: Comparison of DEFENDCLI and Other Methods in Command-Line Detection (Granularity Levels). : Dominant, : Decent, : Moderate, ❍: Basic.
Method OD AC LFD CA AD
Holmes [9] ATT&CK Framework
Unicorn [10] Path Traversal
DepComm [11] Graph Partition
Prographer [12] Graph Partition
Nodlink [13] Steiner Tree Partition
DISTDET [14] Graph Partition
DefendCLI Multi-Level Deviation
1
OD: Restricted Effectiveness in Obfuscation Detection. [15] A significant limitation of many previous approaches lies in their heavy reliance on detecting structural differences or anomalies within execution graphs to identify malicious activities. Although this strategy can be effective for well-defined and previously observed attack patterns, it falls short when addressing the challenges posed by obfuscated command-line execution.
Attackers often leverage obfuscation techniques to disguise malicious behavior, making it difficult for detection systems to identify these threats using traditional anomaly-based or structural methods. The lack of robustness against obfuscation represents a critical vulnerability in these systems, which attackers can exploit to evade detection.
Detecting obfuscated commands is essential for mitigating advanced attack techniques that often rely on subtle modifications to otherwise legitimate processes. However, most existing methods struggle to achieve satisfactory results in this area. For example, Holmes [9] demonstrates decent effectiveness in identifying some obfuscation patterns, but its scope remains limited and fails to provide a comprehensive solution due to its rule-based nature. On the other hand, Unicorn [10], DepComm [11], ProGrapher [12], Nodlink [13], and DISTDET [14] show basic performance in this domain. These methods primarily focus on detecting more straightforward anomalies or structural changes and are not designed to handle sophisticated obfuscation techniques, such as polymorphic commands, encoded payloads, or the use of benign system utilities for malicious purposes. This limitation significantly reduces their effectiveness in combating modern, highly evasive attack strategies, leaving critical gaps in system security.
2
AC: Unclear and Insufficient Attack Correlation. [16] Similarly, these methods often struggle with attack correlation, as identifying relationships between commands typically requires a contextual understanding that goes beyond detecting mere structural deviations. Effective attack correlation is crucial for piecing together the sequence of events within an attack, enabling security teams to understand the full scope and intent of malicious activities.
Reconstructing attack chains through command correlation is vital for analyzing complex threat patterns and attributing malicious activities to specific campaigns. Among existing approaches, Nodlink [13] demonstrates dominant performance by using a steiner tree framework to correlate node-level events. Holmes [9] also has decent success, employing attack chain reconstruction techniques that provide a cohesive view of adversarial behavior. Similarly, Unicorn [10], DepComm [11], and ProGrapher [12] employ graph-based or dependency-oriented methods to trace structural relationships and event dependencies. While effective to a degree, these methods often fall short in capturing the nuanced contextual information necessary to identify more sophisticated and stealthy attack patterns.
In contrast, DISTDET [14] performs basically in this domain due to its lack of robust correlation mechanisms, as it simply deduplicates similar command-line entries without distinguishing between malicious and benign instances. Its limited ability to interpret complex relationships between commands diminishes its effectiveness in reconstructing multi-stage attack chains. This shortcoming is particularly problematic for detecting advanced persistent threats (APTs), which often rely on well-coordinated, multi-layered tactics to achieve their objectives. Overall, the inability of existing methods to fully capture and correlate command-line interactions represents a significant barrier to achieving comprehensive threat detection and analysis.
3
LFD: Incapable of Low-Frequency Events Detection. [17] Furthermore, detecting low-frequency patterns, which are often indicative of novel or highly targeted attacks, remains a critical gap in existing solutions. These patterns typically lack the historical context or recurrence that anomaly-based systems rely upon, making them harder to identify with conventional techniques. Methods that depend solely on anomaly detection tend to produce high false positive rates or miss subtle, context-dependent signals that require a deeper understanding of system behavior.
Detecting rare or low-frequency command-line patterns is particularly vital for identifying emerging threats and sophisticated targeted attacks. However, most existing approaches fail to effectively address this dimension. For instance, Holmes [9] offers moderate rule-based support for low-frequency event detection. Similarly, Unicorn [10], DepComm [11], ProGrapher [12], Nodlink [13], and DISTDET [14] primarily rely on identifying common patterns based on historical profiles of normal activity. However, they overlook the complexities of rare or context-specific anomalies, which frequently evolve and pose significant security challenges. This limitation hampers their ability to detect and respond to advanced stealth attacks that exploit the blind spots of conventional detection mechanisms.
4
CA: Limited Context Awareness. [18] Finally, the lack of context awareness in many existing approaches significantly undermines their ability to differentiate between benign and malicious commands, particularly in dynamic environments where identical commands can have vastly different implications depending on their execution context. Context awareness is critical for accurately interpreting command behavior, as the same command might serve legitimate purposes in one situation but be part of an attack in another. Without the ability to incorporate and analyze contextual data, detection systems often generate high false positive rates or fail to detect subtle malicious activities embedded within otherwise benign operations.
Context awareness enables detection systems to consider factors such as the command’s origin, associated processes, user privileges, and the broader system state, thereby reducing false alarms and enhancing detection precision. Among existing methods, Holmes [9] provides a moderate rule-based context awareness for known attacks, which limits its ability to distinguish between benign and unknown malicious command usage effectively. Similarly, Unicorn [10], ProGrapher [12], Nodlink [13], and DISTDET [14] exhibit basic performance in this area due to minimal integration of contextual data into their detection processes. These methods primarily rely on structural or anomaly-based techniques without considering the broader operational environment in which commands are executed.
In contrast, DepComm [11] shows decent performance using partial contextual information within their method. For instance, the method may take into account some system attributes as dependencies to reduce the attack irrelevance. However, its reliance on context reduction often fails to capture the full spectrum of environmental variables, making them inadequate for detecting sophisticated threats that exploit contextual subtleties. The lack of robust context-aware mechanisms in these approaches leaves significant blind spots in detecting stealthy tactics that mimic legitimate behavior, allowing attackers to exploit contextual ambiguities and bypass traditional detection methods.
III Detection Surface
Figure 3 presents the detection surface of our solution, highlighting its effectiveness compared to previous research, which struggled to achieve accurate detection. DefendCLI focuses on three levels of detection to thoroughly monitor unusual activities, interacting with system audit logs: (1) PC: it identifies unusual system process calls across the OS-level for each endpoint; (2) CL: it analyzes each individual command-line execution within the detected system process chain; (3) NC: it provides information on network connections related to each system process. However, DefendCLI is not a silver bullet to detect all malicious behaviors as part of the advanced security analysis function in EDR. It is often necessary to combine it with other security devices, such as Anti-Virus [19] and NIDS [20], to achieve more comprehensive results. Our solution focuses primarily on attack activities that can be observed and monitored from system audit logs, which provide detailed behavioral and operational information.
1
PC: Unusual Interacted System Process Calls. Observing system process calls is crucial for detecting abnormal OS activities, as legitimate software typically does not invoke unrelated applications or scripts. Despite attackers embedding malicious code in trusted software, a technique known as Living off the Land Attack [21], process chains still provide transparency by exposing deviations from normal behavior [22]. For instance, when malware is injected into a legitimate process, it spawns a new thread within the same namespace, altering system memory allocation. If a thread attempts to access another process’s memory, the system audit log records it as interprocess access, making such activities detectable through process call interactions.
Figure 3: Detection Surface: Given two similar attack chains, previous detection methods failed to distinguish the structural differences, leading to an inability to correctly identify the mimicry attack (right) resulted in blind classification.
2
CL: Suspicious Command-Line Executions. The effectiveness of system process chains is often compromised when malware embeds itself within legitimate processes, reducing visibility and detectability. Our solution addresses this by analyzing command-line attributes of flagged system processes within a process chain. This approach, which inspects command-line inputs from both interactive shells and OS API executions, enhances security analysts’ ability to swiftly review EDR results when process chain investigations become challenging. Certain command-line interactions are critical for OS functions and attack phases, making them unavoidable despite evasion tactics.
For example, the Turla APT Group [23] employs malware that leverages .NET or C++ to communicate with C2 servers, concealing malicious command-line activity. However, system process call analysis can still reveal anomalies, such as, Explorer.exe $\rightarrow$ Powershell $\rightarrow$ csc.exe (command-line: /noconfig…4krwc2ua.cmdline) $\rightarrow$ WMI.exe (command-line: tasklist). Notably, csc.exe rarely invokes 4krwc2ua.cmdline from the .NET library and should not trigger WMI.exe to execute tasklist, marking it as a potential anomaly.
3
NC: Rarely Seen External Network Connections. To boost detection capabilities, we integrate basic network data along with process-level and command-line-level analysis. Unlike network intrusion detection that uses detailed network attributes from packet inspection, the system audit log only records the essential network information IP: Port. The advantage of this method is its ability to link network connections directly with specific system processes. In the context of the Trula APT case, the detection outcome will be supplemented with network connections such as, Explorer.exe $\rightarrow$, Powershell $\rightarrow$, csc.exe (command-line: /noconfig/…//4krwc2ua.cmdline) $\rightarrow$, WMI.exe (command-line: tasklist) $\rightarrow$, and connection: 154.26.156[.]62:4444. This series reveals an unusual external connection by csc.exe, a fundamental component of .NET, to execute system commands through the 4krwc2ua.cmdline library object.
Figure 4: Attack-Clause Sketch and Attack-Evidence Awareness: The attack-clause sketch module includes (1) provenance construction for building behavior graphs and (2) command-line evaluation to refine weights using known attacks. The attack-evidence awareness module handles (3) anomaly detection for unknown threats and (4) reporting via RAG-LLM.
IV DefendCLI
To address the above challenges, we propose DefendCLI, a comprehensive system built with significant engineering effort to support actionable threat verification. We begin by outlining the system architecture, then describe each component following its execution workflow. For each part, we explain the design rationale and discuss the techniques used to implement our solution.
IV-A System Overview
Figure 4 illustrates the architecture of the DefendCLI system, which adopts a highly modular design to facilitate extensibility and clarity. The system is composed of two core components: Attack-Clause Sketch, responsible for constructing the attack provenance graph, and Attack-Evidence Awareness, dedicated to threat detection.
1
Attack-Clause Sketch. The attack-clause sketch can be constructed to systematically represent and trace the logical structure of an attack as discussed below.
IV-A1 Attack Provenance Construction
It begins with the attack provenance construction module, which is designed to capture the interactive relationships among multi-level system events. This module is carried out in three key steps: (1) attack-irrelevance reduction: a reduction method is used to simplify the graph while maintaining its critical structural and relational integrity. (2) attack-chain causal inference: a causal inference technique is applied to identify potential attack-related edges and nodes through various weighting schemes, enhancing the graph’s interpretability and analytical value. (3) interphase attack association: a solution is implemented to assess potential relationships between partitioned subgraphs, ensuring that multi-staged attacks are effectively correlated.
IV-A2 Command-Line Differentiation Evaluation Based on Known Facts
Following the attack provenance construction phase, the system transitions to the evaluation of command-line actions. This phase integrates multiple evaluation strategies. Initially, process chains and their associated command-line executions are analyzed using Sigma rules [24]. These rules, crafted by security experts and organizations, identify specific malicious signatures, significantly reducing false positives when malicious actions match these signatures. Next, undetected command-line actions are cross-referenced against a repository of defense evasion rules [25, 26], which address common system command-line patterns often exploited for evasion. The evaluation results are then used to adjust the weights within the attack provenance graph, improving the detection accuracy.
2
Attack-Evidence Awareness. The attack-evidence awareness module is detailed below, highlighting the core elements that enable effective identification and contextualization of attack-related evidence.
IV-A3 Explainable Anomaly Detection for Unknown Attacks
As attack clauses accumulate, the corresponding evidence becomes increasingly explicit and compelling, necessitating timely detection. In the explainable anomaly detection phase, the system conducts a thorough analysis of the partitioned attack provenance graph. This process begins with Information Path (InfoPath) traversal, which identifies and ranks anomalous paths based on weighted scoring. While effective for known threats, rule-based detection alone is insufficient to capture previously unseen attack behaviors. To address this limitation, we propose a SimHash-based similarity framework that analyzes unknown command-line instances. This framework informs the selection of embedding strategies that maximize intra-class separability, thereby improving contextual awareness of novel attack patterns. To further enhance robustness, we incorporate an unsupervised voting mechanism that aggregates outputs from multiple machine learning models, reducing individual model bias and improving detection stability. This integrated approach improves anomaly detection performance while maintaining interpretability. Additionally, adaptive hyperparameter tuning is employed to optimize system parameters in real time, ensuring consistent performance across diverse and dynamic operational contexts.
IV-A4 Report Generation and Recommendation
To enhance SecDevOps efficiency, we employ the RAG-GPT4o model to automate the generation of comprehensive analysis reports derived from anomaly detection outputs. This model systematically ranks snapshot-based recommendations, highlighting the Top-5 most suspicious activities with contextual insights. These actionable insights include precise interpretations of suspicious command-line activities, inferred adversarial tactics, and potential threat indicators. By summarizing critical anomalies concisely and contextually, security analysts can quickly evaluate whether a host is compromised, significantly reducing cognitive overload and mitigating alarm fatigue. This streamlined approach empowers teams to swiftly prioritize responses, enhancing overall threat management capabilities
Figure 5: Comparison of (a) Information-Flow Focused Structure and (b) Refined Command-Line Focused Structure: Both network and command-line details are stored as memory attributes within process nodes. The refined structure omits file nodes to reduce redundancy and lower false alarm rates, as files are rarely the root cause [27].
IV-B Attack-Clause Sketch
The attack-clause module comprises two key components: (1) attack provenance construction and (2) command-line differentiation evaluation based on known attack patterns. These phases are critical for correlating attack clauses and quantifying the associated risk levels.
1
Attack Provenance Construction. To address the limitations of detecting malicious command-line activities, we design a refined attack provenance structure. This enhanced representation treats command-line inputs not merely as deviations within the global behavior graph but also as standalone semantic features. By incorporating command-line semantics as independent signals, the structure improves the granularity of behavioral differentiation and enhances the detection of subtle or previously unseen attacks. Figure 5 provides a comparative visual representation of the improvements in constructing the provenance graph.
IV-B1 Attack Provenance Structure
Figure 5(a) illustrates the conventional approach, termed the Information-Flow Focused Structure, as described in [28]. This method employs a heterogeneous graph composed of diverse node types, including reverse shell processes, system processes, command-line inputs, file paths, and network activities.
In contrast, Figure 5(b) depicts our proposed Refined Command-Line Focused Structure. Unlike the heterogeneous architecture, our design adopts an isomorphic graph structure with uniform node types, centering on system processes. Command-line and network information are encapsulated as memory attributes within each process node, reducing structural noise while preserving semantic context.
This refined design yields multiple advantages. By unifying node types, it streamlines graph representation and computation, enabling more effective application of graph-based analytics such as embedding techniques, anomaly detection algorithms, and structural similarity metrics. The reduced complexity and consistent node attributes enhance the system’s ability to detect and differentiate subtle anomalies in attack behavior.
IV-B2 Attack Irrelevance Reduction
To improve the clarity and relevance of attack provenance analysis, we introduce a reduction technique that eliminates non-correlated events from the provenance graph. This targeted reduction is applied at both the node and edge levels, mitigating the dependency explosion problem, a well-known challenge in provenance analysis [29].
At the node level, we remove subgraphs that do not contain any command-line elements. This decision is based on two key observations. First, normal system behavior seldom involves extensive command-line invocation. Second, effective incident investigation often requires the presence of command-line clauses to trace adversarial intent. By isolating command-line-relevant components, the proposed reduction method streamlines the graph, reduces analytical overhead, and minimizes the likelihood of misleading dependencies.
At the edge level, a critical concern is the presence of loops and cycles within the provenance graph, which can compromise structural integrity and impede accurate data correlation. To address this issue, we transform the graph into a Directed Acyclic Graph (DAG) [30]. This transformation involves systematically identifying and eliminating cyclic dependencies by removing looped edges, thereby ensuring acyclic structure and facilitating more reliable provenance analysis.
Figure 6: Attack-chain causal inference: the process begins by calculating node weights using PageRank and Shortest-Path Betweenness Centrality. Subsequently, these node weights are used to determine the edge weights.
Figure 7: Interphase attack association: a Leiden community detection method is used to correlate interphase attacks.
IV-B3 Attack-Chain Causal Inference
Identifying deviations that align with plausible attack inferences remains one of the most challenging problems in attack provenance analysis. To address this, we adopt explainable and practical techniques to isolate abnormal instances embedded within the inference process. Figure 6 illustrates our approach for identifying critical elements within the complete provenance graph, enabling focused and interpretable anomaly detection.
$\blacksquare$ How can we identify the most unusual elements traversed through reachable paths? Attack-relevant elements are often rare and obscured by a large volume of benign activities. To surface these critical nodes, we leverage the PageRank algorithm [31], originally developed to rank web pages by assessing their relative importance based on link structures. In this context, PageRank models the provenance graph as a network, where an edge from node $A$ to node $B$ is interpreted as a “vote” indicating the significance of $B$. Nodes with fewer incoming connections are considered more anomalous, reflecting their rarity within the observed activity flow.
Mathematically, the PageRank ($PR$) of a node is computed by summing the $PR$ values of all nodes linking to it, each divided by the number of outgoing links from the source node, and then scaled by a damping factor. This process is defined as:
$PR(v)=(1-d)+d\sum_{u\in\text{In}(v)}\frac{PR(u)}{L(u)},$
where $d$ is the damping factor (typically set to 0.85), $\text{In}(v)$ is the set of nodes linking to node $v$, and $L(u)$ is the number of outgoing links from node $u$. The algorithm operates iteratively, recalculating scores across the graph until convergence.
Unlike its traditional application, we normalize the final $PR$ scores to reflect rareness rather than prominence. This inversion highlights nodes that are less connected and potentially anomalous, enabling the effective identification of suspicious or atypical attack elements within the provenance graph.
$\blacksquare$ How can we identify the root cause of an attack among all accessible elements? [32] Identifying the root cause is essential for understanding attack progression. However, we do not define the root cause merely as the origin point in a source-to-destination path. Instead, we conceptualize it as a pivotal bridge node that connects early-stage activities to subsequent malicious behaviors.
To uncover such bridge nodes, we employ the shortest-path betweenness centrality algorithm [32], which quantifies the importance of a node based on the number of shortest paths that traverse it. Specifically, this measure captures how frequently a node serves as an intermediary on the shortest paths between all pairs of nodes in the graph. In the context of attack provenance, such nodes often correspond to compromised services or applications [33] that enable lateral movement or privilege escalation.
The computation involves the following steps: (1) Identify all node pairs $(s,t)$, where $s$ is a source and $t$ a destination. (2) For each pair, compute all shortest paths between $s$ and $t$. (3) For each path, count the occurrences of node $v$ along the path, excluding cases where $v=s$ or $v=t$. (4) Aggregate these counts across all pairs to obtain the betweenness centrality of node $v$.
This metric enables the identification of structurally significant nodes that are critical to the attack’s progression, thereby supporting effective root cause analysis in complex networked environments.
Figure 8: Explainable anomaly detection mechanism for unknown attacks: Ranked InfoPath will be used for the anomaly detection. SimHash recommender will examine the difference in command-line executions. Finally, the Police-Witness Voting Policy will select the most anomalous InfoPath for evaluation.
Figure 9: Example of SimHashed command-line executions: the red box of $\blacksquare$ is malicious.
$\blacksquare$ How can we integrate both causal inference methods for enhanced analysis? Both PageRank and shortest-path betweenness centrality evaluate the significance of nodes within the provenance graph, yielding node-level importance scores. To leverage the strengths of both methods within the detection framework, we transform node weights into edge weights, an essential step for enabling InfoPaths traversal during anomaly detection.
Given a pair of nodes $m$ and $n$ connected by an edge, we compute the corresponding Edge Weight ($EW_{(m,n)}$) as the sum of their individual node scores:
$EW_{(m,n)}=PR_{(m)}+PR_{(n)}+CB_{(m)}+CB_{(n)},$
where $PR$ denotes the PageRank score and $CB$ represents the shortest-path betweenness centrality score of a node.
To ensure comparability and stability across the graph, all edge weights are subsequently normalized to the range $[0,1]$. This normalization facilitates consistent interpretation and supports a unified, edge-centric approach to anomaly detection within the attack provenance framework.
IV-B4 Interphase Attack Association
Correlating multi-stage attacks within a single group poses a significant challenge, as it requires a holistic understanding of the full attack chain. Establishing meaningful associations between disparate phases demands both sufficient and reliable contextual information to uncover subtle dependencies across stages.
To address this challenge, we adopt a graph partitioning strategy using community detection to group strongly connected and semantically related nodes while isolating unrelated components. This is commonly achieved via the Louvain algorithm [34], a modularity optimization method frequently applied in attack graph analysis [35]. Louvain iteratively optimizes modularity until a stable partitioning is reached.
However, a well-documented limitation of Louvain [36] is its tendency to produce arbitrarily weakly connected communities. In extreme cases, the resulting communities may become disjoint after successive iterations, leading to inaccurate or incomplete data extraction, as reported in Depcomm [11]. Such fragmentation impairs the reliability of inter-phase correlation.
To mitigate this issue, DefendCLI replaces Louvain with the Leiden algorithm [37] for community detection, as illustrated in Figure 7. Empirical studies [38] have shown that Leiden not only improves partition quality by preserving community cohesion but also enhances computational efficiency, enabling faster analysis of large-scale graphs.
Furthermore, each detected community is assigned a community score, which is subsequently incorporated into the InfoPaths search process to reinforce the correlation of multi-stage attack behaviors across the provenance graph.
2
Command-line Differentiation Evaluation (Known Facts). Based on real-world feedback, it is inadvisable to disregard expert knowledge during early-stage detection.
$\blacksquare$ How can we detect known attack behaviors? To support practical deployment, we evaluate detected system process chains using Sigma rules [24] via SIEM platforms (e.g., Elasticsearch). If a matching pattern is identified, the process chain is labeled as malicious and assigned a High-Risk Impact with an associated Risk Score ($RS$).
$\blacksquare$ How can we detect known defense evasion techniques? Process chains are further analyzed using a curated set of evasion rules [25, 26]. These rules adjust the weights in the attack provenance graph based on rule confidence levels. Evasion indicators are categorized into Medium-Risk and Low-Risk classes, each assigned a corresponding $RS$.
$\blacksquare$ How can we integrate both methods into attack provenance analysis? Following command-line evaluation, identified patterns are mapped to the provenance graph using regular expressions [39]. For each matching edge between nodes $m$ and $n$, the Refined Edge Weight ($ReEW_{(m,n)}$) is computed as:
$ReEW_{(m,n)}=RS\times EW_{(m,n)},$
where $EW_{(m,n)}$ is the original edge weight. Updated edges adopt the new weight, while unmatched nodes and edges retain their original values.
IV-C Attack-Evidence Awareness
This module comprises two primary components: (1) explainable anomaly detection to identify unknown attacks and (2) automated threat report generation with actionable insights. By analyzing accumulated attack clauses, these components work in tandem to enhance understanding of attack patterns and support effective threat identification.
3
Explainable Anomaly Detection (Unknown Attacks). To detect novel attacks based on accumulated clauses, we design an explainable anomaly detection framework, as illustrated in Figure 8.
IV-C1 Information Retrieval (InfoPath Search)
A central challenge in attack provenance analysis is determining how to retrieve malicious information from large volumes of provenance data using attack clauses. To this end, we propose InfoPath Search, a retrieval method designed to efficiently extract relevant attack-related information.
$\blacksquare$ How can we retrieve malicious information based on attack clauses? InfoPaths are derived from each partitioned cluster based on interphase attack associations. We employ the Dijkstra algorithm [40] to compute the shortest paths between node pairs.
Traversal begins from nodes with zero in-degree and terminates at nodes with zero out-degree, when applicable; otherwise, all paths are considered. This approach ensures that, in a directed provenance graph, the shortest InfoPaths reflect plausible causal attack chains from source to sink.
$\blacksquare$ How does information retrieval relate to causal inference and command-line evaluation? In our framework, higher ($ReEW$) denote higher significance in causal inference. However, since shortest-path algorithms prioritize lower weights, we invert the edge weights using $\frac{1}{ReEW}$ to ensure that rare and anomalous paths are prioritized. This transformation aligns the retrieval process with the objective of uncovering unusual and potentially malicious patterns.
$\blacksquare$ How does information retrieval support interphase attack correlation? Shorter InfoPaths typically indicate more suspicious activity. To enhance interphase attack correlation, path lengths are further adjusted using the community score (CS), favoring paths that traverse high-impact communities, that is defined as $\frac{1}{ReEW\times CS}$. The final InfoPath rankings are based on two factors: (i) effective path length and (ii) the diversity of system processes involved. This accounts for the practical observations that malware often leverages various system processes to evade detection [41].
IV-C2 Command-Line Differentiation Evaluation Based on Unknown Facts (SimHash Method)
To improve the identification of novel attacks, we propose a framework that delivers an impartial anomaly detection system. This framework bridges the gap between algorithmic uncertainty and the capability to detect previously unseen command-line activities.
$\blacksquare$ How can we detect subtle context deviations to effectively address complex multi-stage attacks? Previous research often overlooks the impact of varying embedding techniques on security datasets, leading to reduced sensitivity across diverse data types. This oversight diminishes the granularity of detection and restricts context awareness to minor deviations, an essential limitation when addressing novel and sophisticated attacks.
To address this issue, we utilize the SimHash [5] algorithm to randomly assess similarity within post-embedding command-line data. Our approach incorporates embedding techniques such as Word2Vec [42], Doc2Vec [43], and FastText [44], which have been employed in previous studies. This integration aims to mitigate the impact of vectorization on different types of security audit data.
SimHash generates a unique hash value for each embedded command-line entry, as shown in Figure 9. The process begins by transforming command-line data into vector representations using Word2Vec [42], Doc2Vec [43], and FastText [44], which capture semantic relationships and contextual dependencies within command sequences.
Then, a weighted hashing function is applied to generate a fixed-length binary signature for each embedded command-line entry. This technique ensures that similar command-line sequences produce hash values with minimal Hamming distance, whereas significantly different sequences result in more distinct hashes. By computing and analyzing these distances, the approach quantifies the similarity between different command-line entries.
Finally, the embedding method that best distinguishes the differences among all evaluated command-line inputs based on SimHash similarity scores is selected for anomaly detection. This process ensures that vectorization effectively preserves critical distinctions within security audit data, enhancing the detection of novel attacks while reducing unnecessary false positives.
$\blacksquare$ How can we objectively detect anomalies based on command-line behavioral differentiation? Recent research trends favor the use of neural networks for anomaly detection. However, these methods often require large volumes of high-quality training data and substantial computing resources. In contrast, we believe that a cost-effective approach is well-suited for attack provenance examination, which typically demands lightweight on-site deployment. To achieve this goal, we developed a multifactor-based detection mechanism inspired by criminal investigation methods.
Various classifiers (IsolationForest [45], LocalOutlierFactor [46], OneClassSVM [47], EllipticEnvelope [48], KMeans [49], and GaussianMixture [50]) then respond based on their respective algorithms, ensuring diverse analytical perspectives. This approach mitigates algorithmic bias by preventing over-reliance on similar features. The final decision is determined through a binary voting system [51], where the suspect receiving the most votes is identified.
In this context, suspects are represented by InfoPaths embedded with command-line vectors, while multiple classifiers provide supporting evidence, culminating in the voting process. Ultimately, InfoPaths identified as malicious and exhibiting significant command-line differentiation are selected for further analysis.
TABLE II: Effect of Hyperparameters on Adaptive System
Hyperparameters Positive Effect Negative Effect
$C$ $\uparrow C\uparrow$ DR $\uparrow C\uparrow$ FAR
RSRS $\uparrow RS\downarrow$ FAR $\uparrow RS\downarrow$ DR
NNNN $\uparrow NN\uparrow$ DR $\uparrow NN\uparrow$ FAR
VSVS $\uparrow VS\uparrow$ DR $\uparrow VS\uparrow$ FAR
IV-C3 Adaptive Interface for Dynamic Environments
Machine learning-based detection mechanisms often demonstrate limited effectiveness in real-world environments due to the dynamic nature of software and hardware ecosystems. Frequent updates and evolving system behavior can degrade model performance over time, necessitating retraining and continuous evaluation, both of which introduce significant human and computational overhead.
To mitigate these challenges, we propose an adaptive interface that allows external users to adjust key hyperparameters that influence False Alarm Rate (FAR) and Detection Rate (DR), with particular emphasis on adjusting classification thresholds. This user-configurable approach reduces the opacity of AI-based systems and supports improved operational alignment with diverse deployment conditions.
Through extensive experiments, we identify several hyperparameters that significantly impact system performance. These are summarized in Table II, and can be tuned according to local environmental characteristics. Automated tuning, while convenient, often leads to increased false alarms in heterogeneous settings, reinforcing the need for manual control.
$\blacksquare$ Influence of Risk Score (RS) based on known attack facts: As described in Section IV-B4, expert knowledge is incorporated through command-line differentiation, categorizing risk into three levels: High (H), Medium (M), and Low (L), each associated with a corresponding RS. These default values are set as: {H = 2.5, M = 2.0, L = 1.5}, reflecting the confidence level in known attack indicators.
$\blacksquare$ Influence of Contamination ($C$) based on audit data quality: Real-world audit data often contain inconsistencies, making it difficult to accurately distinguish between benign and malicious activity. To account for data imbalance, we introduce the Contamination parameter $C$, which is integrated into supported classifiers to control for class skew. The default value $C=0.5$ represents a moderate assumption of data imbalance.
$\blacksquare$ Influence of Number of Neighbors (NN) and Vector Size (VS): Data intensity varies between environments, which requires flexibility in classifier configuration. The Number of Neighbors ($NN$) and Vector Size ($VS$) parameters help scale detection strength. Lower values are recommended for smaller datasets, while higher values are preferable in high-volume environments. The default settings are $NN=300$ and $VS=60$.
4
Report Generation and Recommendation. Alarm fatigue, a widespread problem in cybersecurity, occurs when security analysts are inundated with excessive alerts, resulting in missed critical threats and delayed responses. To mitigate this challenge, we introduce a snapshot alarm mechanism powered by a GPT model enhanced through a Retrieval-Augmented Generation (RAG) framework based on GPT-4o. The RAG system incorporates a curated set of trusted security reports [52, 53, 54], significantly improving the model’s ability to detect, contextualize, and prioritize emerging threats.
Figure 10 illustrates the GPT model recommendation workflow. The mechanism ranks all suspicious InfoPaths and highlights the Top-5 for further inspection, allowing analysts to focus on the most critical alerts and enhancing situational awareness prior to full investigation. In addition, the model conducts an in-depth analysis of detected command-line activity, offering clear explanations of intent and potential impact. This contextual intelligence supports more accurate threat identification, reduces false positives, and streamlines decision-making for security operations teams.
Figure 10: GPT-Powered Snapshot Alarm and Threat Prioritization: An overview of the GPT model’s recommendation process, which ranks suspicious InfoPaths and provides in-depth analysis of command-line activities.
V Evaluation
The evaluation is divided into overall evaluation and ablation evaluation. The overall evaluation compares our approach with state-of-the-art methods using precision, recall, and $F_{1}$-score on selected academic and industrial attack datasets. The ablation evaluation conducts a detailed functional analysis to assess the effectiveness of each component in our solution, providing deeper insights into its impact and performance on discussed RQs and limitations.
V-A Attack Datasets
V-A1 DARPA Engagement Series Datasets (Linux and Unix-Like Servers)
The DARPA Transparent Computing Engagement Series (TCES) [55] are designed to support cybersecurity research by providing real-world cyber attack traces, system logs, and network activity. Due to most of related work are evaluated on the Engagement-3 (E3) dataset, to make a fair comparison, we also use the E3 dataset as the benchmark. The E3 dataset captures multi-user cyber attack scenarios with a mix of legitimate and malicious activities, including privilege escalation, lateral movement, and data exfiltration, making it ideal for anomaly detection and forensic analysis. It includes host-level logs, file system events, system calls, and process activities, making them valuable for developing machine learning models, behavioral analytics, and proactive cyber defense strategies. Based on the attack scenarios, the E3 dataset is divided into E3-Trace, E3-THEIA and E3-CADETS.
TABLE III: GitHub Availability, Reproducibility and Real-Time Detection of Related Work
Method A RP RT C
Holmes [9] S&P
Unicorn [10] NDSS
DepComm [11] S&P
Prographer [12] Usenix Security
Nodlink [13] NDSS
DISTDET [14] Usenix Security
TABLE IV: Licensed Commercial EDR Solutions
Tool AV WS CTI RT
Microsoft Defender [56]
Symantec EDR [57]
Karspersky EDR [58, 59]
TABLE V: Node-Level Results on DARPA E3 Datasets
Method E3-Trace E3-THEIA E3-CADETS
Precision Recall $F_{1}$ Score Precision Recall $F_{1}$ Score Precision Recall $F_{1}$ Score
Holmes [9] $1.35\times 10^{-3}$ 0.74 $2\times 10^{-3}$ $3.61\times 10^{-3}$ 0.98 0.01 $2.84\times 10^{-3}$ 0.95 0.01
Unicorn [10] $3.20\times 10^{-5}$ N/A N/A $1.86\times 10^{-4}$ N/A N/A $1.25\times 10^{-4}$ N/A N/A
DepComm [11] $2.3\times 10^{-3}$ 1 $4.6\times 10^{-3}$ $1.2\times 10^{-2}$ 1 0.02 $1.2\times 10^{-2}$ 1 0.02
Nodlink [13] 0.25 0.98 0.40 0.23 1 0.37 0.14 1 0.25
DefendCLI 0.41 1 0.58 0.38 1 0.55 0.2 1 0.33
TABLE VI: Attack Scenario Checklist (E3)
Dataset Attack Scenario TTPs
TRACE Firefox Exploit via Ads Drive-by Execution
Malicious Extension Drops Payload DLL Injection
Phishing Email Installs Backdoor Credential Theft
THEIA Firefox Exploit Injects Script Remote Execution
Fake Extension Steals Passwords Keylogging Attack
CADETS Web Server Exploited Remotely Process Injection
Fake Update Installs Malware Persistence Attack
Lateral Movement via Credentials Privilege Escalation
(a) Attack-Related Command-Line (E3-TRACE)
(b) Attack-Related Command-Line (E3-THEIA)
(c) Attack-Related Command-Line (E3-CADETS)
(d) InfoPath-Level Result (E3-TRACE)
(e) InfoPath-Level Result (E3-THEIA)
(f) InfoPath-Level Result (E3-CADETS)
Figure 11: Comparison of attack-related command-line and InfoPath-Level results on E3 datasets.
V-A2 Industrial Real-Time Detection (Microsoft Windows End-User and Servers)
The Industrial Real-Time Detection dataset [60] is designed to support cybersecurity research by providing real-world, real-time telemetry from virtual machines (VMs) hosted on Microsoft Azure, enabling the evaluation of real-time detection and response capabilities for defense solutions. The dataset captures system logs, network activity, and security events in an industrial environment, simulating active cyber threats and operational challenges as they unfold. By including both benign and malicious activities, it allows researchers to study and enhance real-time threat detection, anomaly identification, and rapid intrusion response in cloud-based infrastructures. With realistic attack scenarios such as unauthorized access, privilege escalation, and lateral movement, this dataset is particularly valuable for advancing machine learning-driven security solutions, continuous monitoring systems, and proactive cyber defense strategies, emphasizing the importance of real-time detection and mitigation in industrial cloud environments.
TABLE VII: Attack-TTPs and Node-Level Results on Industrial Real-Time Detection
Method Attack Scenario #1 (A1) Attack Scenario #2 (A2) Attack Scenario #3 (A3)
Precision Recall $F_{1}$ Score Precision Recall $F_{1}$ Score Precision Recall $F_{1}$ Score
Microsoft Defender [56] 0.04 1 0.08 0 1 0 0 1 0
Symantec EDR [57] block block block 0.13 1 0.23 0.04 1 0.08
Kasparsky EDR [58] block block block block block blcok 0.09 1 0.10
Nodlink [13] (TTPs) 1 1 1 1 1 1 1 1 1
Nodlink [13] (Node-Level) 0.04 1 0.08 0.19 1 0.32 0.05 1 0.10
DefendCLI (TTPs) 1 1 1 1 1 1 1 1 1
DefendCLI (Node-Level) 0.08 1 0.15 0.38 1 0.55 0.18 1 0.31
TABLE VIII: Attack Scenario Checklist (Industrial)
Scenario APT TTPs
1 Silent Lynx [61, 62] PasteBin Multi-Staged Powershell
2 DriftingCloud [63, 64] MITM PHP Chopper Webshell
3 Evilnum [65] Spear Phishing Python RAT
V-B GitHub Availability, Reproducibility and Real-Time Detection
Table III shows the availability (A), reproducibility (RP), and real-time detection capability (RT) of related work based on big-4 security conferences (C). The research work of DISTDET [14] and Prographer [12] are not publicly available on GitHub or other sources. Therefore, our comparative study focuses on the remaining work. The verification of availability was conducted on 7 February 2025.
V-C Licensed Commercial EDR Solutions
Table IV presents the detection capabilities of commercial EDR solutions across four key security features: anti-virus (AV), web security (WS), cyber threat intelligence (CTI), and real-time (RT) protection. For our evaluation, we obtained commercial licenses for Microsoft Defender [56], Symantec EDR [57], and Kaspersky EDR [58, 59]. Each EDR solution was updated to the latest version, ensuring the most recent threat database was used during testing.
V-D Attack Provenance Evaluation Criterion
We use a multi-level criterion to evaluate our work, which is introduced as follows:
  • InfoPath-Level: The InfoPath-level evaluation criterion was introduced by DefendCLI, which detects anomalous attack chains by analyzing multiple paths between source and destination nodes. InfoPaths represent the interactions between system behaviors and serve as a core metric in both real-time industrial detection scenarios and our ablation studies.
  • Node-Level The node-level criterion, first introduced in Nodlink [13], addresses the limitations of prior approaches that evaluated entire attack provenance graphs often achieving perfect precision while inadvertently including many non-attack-related events. Instead, this criterion focuses on identifying and quantifying attack-relevant nodes within InfoPaths, offering a more granular and accurate reflection of actual attack activity.
  • TTPs-Level The TTPs-level criterion is used to compare detection capabilities against commercial EDR tools. Since many EDRs terminate execution upon detecting an attack, limiting further behavior observation, we assess the number of detected activities associated with adversarial tactics, techniques, and procedures (TTPs). This allows for a direct performance comparison between our solution and commercial detection tools.
V-E Overall Evaluation
V-E1 Detection Results on DARPA E3 Benchmark
Table VI presents the threat model of the three datasets in the E3 benchmark, while Table V compares node-level detection results with state-of-the-art research. Our analysis shows that while all existing methods can detect E3 attack scenarios, their node-level results vary significantly. This discrepancy highlights the presence of substantial non-attack-related information in their alarms, leading to lower detection precision and reduced $F_{1}$ scores.
Our solution DefendCLI, improves detection by focusing on command-line-level analysis, achieving a precision improvement of 1.6 $\times$ over the best existing method Nodlink [13]. Figure 11 presents the results at the InfoPath-level and the command-line activities related to the detected attack identified by DefendCLI. This improvement demonstrates that DefendCLI provides more concise results, detecting E3 attack scenarios with fewer nodes while delivering more actionable insights for threat incident verification, significantly enhancing detection efficiency.
(a) Attack-Related Command-Line (Real-Time)
(b) Real-Time InfoPath Alarm (Attack Scenario #1)
(c) Real-Time InfoPath Alarm (Attack Scenario #2)
(d) Real-Time InfoPath Alarm (Attack Scenario #3)
Figure 12: Comparison of attack-related command-line and InfoPath alarms in real-time detection.
V-E2 Detection Results in the Industrial Real-Time Detection
Table VIII presents the threat model for industrial real-time detection, incorporating modern APT threats relevant to 2025. These advanced threats are designed to evade modern defense strategies, enabling a more comprehensive evaluation of our solution’s detection capabilities in complex industrial environments.
[…]
Figure 12 illustrates the attack-related command-line activities and the number of real-time InfoPath alarms identified by DefendCLI. The industrial environment presents greater complexity than the E3 datasets due to a significant number of active users generating normal activities during evaluation. Despite this, our results show that DefendCLI maintains a sound alarm rate relative to the volume of generated events, demonstrating its practicality for real-world deployment.
(a) Attack-Chain Causal Inference (E3-TRACE)
(b) Attack-Chain Causal Inference (E3-THEIA)
(c) Attack-Chain Causal Inference (E3-CADETS)
(d) Attack-Chain Causal Inference (Attack Scenario #1)
(e) Attack-Chain Causal Inference (Attack Scenario #2)
(f) Attack-Chain Causal Inference (Attack Scenario #3)
Figure 13: Comparison of attack-chain causal inference based on the number of weighted edge count versus the number of attack-related weighted edge count.
V-F Ablation Evaluation
V-F1 Evaluating the Interoperability of Attack Provenance Examination Through Attack-Clause Sketch and Evidence Awareness
Our solution incorporates a variety of sophisticated algorithms to enhance the detection capabilities of attack provenance analysis. In this section, we present a detailed assessment, breaking down the contribution of each module to demonstrate their individual effectiveness within our framework.
$\blacksquare$ How well does attack-chain causal inference identify unusual elements within the attack provenance graph? The effectiveness of attack-chain causal inference in detecting unusual elements within the attack provenance graph is evaluated based on the comparison between the total number of weighted edges and the subset of edges directly associated with attack activities. Figure 13 illustrates this comparison in different scenarios, highlighting the ability of our method to isolate suspicious elements from normal system interactions.
The results demonstrate that our causal inference model successfully assigns precise and meaningful weights to attack-related edges, ensuring that critical attack components are effectively highlighted while enabling efficient path traversal. This mechanism allows for a clear distinction between malicious activities and benign system interactions, minimizing noise, and improving detection accuracy. The observed disparity between the total number of weighted edges and the subset associated with attacks further validates the model’s ability to prioritize attack-relevant edges, effectively filtering out irrelevant connections and improving the interpretability of the attack provenance graph.
TABLE IX: The Use and Not Use of Community Detection and SimHash Recommender on Node-Level $F_{1}$ Score
Dataset Community Detection SimHash (Blue)
Y N W2V D2V FT
E3-TRACE 0.58 0.47 0.58 0.48 0.50
E3-THEIA 0.55 0.54 0.31 0.42 0.55
E3-CADETS 0.33 0.30 0.33 0.31 0.32
A1 0.15 0.13 0.14 0.14 0.15
A2 0.55 0.31 0.48 0.30 0.55
A3 0.31 0.31 0.31 0.18 0.26
$\blacksquare$ How successfully does interphase attack association enhance the correlation of multi-staged attacks? Figure 14(a) compares the use of community detection in analyzing the provenance of the attack, where Y/N denotes whether community detection is applied. The blue result highlights the better outcome. We reconstructed the detected InfoPaths as weakly connected attack provenance subgraphs to assess how many reachable elements could be accurately correlated.
For the E3-TRACE and E3-THEIA datasets, we observed that without community detection, our solution results in a higher false alarm rate (FAR) while attempting to identify all attack activities. In contrast, for the E3-CADETS dataset, particularly in industrial attack scenarios #1 and #2, we found that while the absence of community detection may lead to fewer results, some attack activities go undetected. However, for industrial attack scenario #3, we observed no significant difference, as all elements were reachable regardless of the use of community detection.
Table IX shows the node-level $F_{1}$ score with or without the use of community detection. The experimental results can support our claim that the application of community detection improves correlation accuracy for identifying interphase attacks, leading to more precise and effective threat analysis.
$\blacksquare$ How proficient is the SimHash recommender in distinguishing unknown command-line activities? Figure 14(b) compares the use of the SimHash recommender in analyzing the provenance of the attack, where W2V, D2V, and FT represent the use of Word2Vec, Doc2Vec, and FastText, respectively. The blue result highlights the selection of the SimHash recommender. For the E3-TRACE, E3-CADETS and attack scenario #3, SimHash recommender selected Word2Vec as the method of vectorization because it has the maximum score to distinguish the command-line differentiation. On the other hand, for the E3-THEIA, attack scenario #1 and #2, SimHash recommender selected FastText as the method of vectorization based on the same standard.
Similarly, we observed that for E3-TRACE and attack scenarios #2 and #3, SimHash generated fewer results while still detecting all attack behaviors, effectively reducing the FAR. In contrast, for E3-THEIA, E3-CADETS, and attack scenario #1, SimHash produced a larger set of results, ensuring comprehensive attack detection. However, other methods, despite yielding fewer results, failed to identify several malicious behaviors.
Table IX presents the node-level $F_{1}$ score with and without the use of SimHash recommender. The results demonstrate that incorporating the SimHash recommender enhances context awareness by effectively differentiating command-line activities, leading to more precise threat analysis.
(a) Community Detection
(b) SimHash Recommender
Figure 14: Comparison of the use with or without community detection and SimHash recommender based on the count of InfoPaths detected.
(a) Diverse OS Platforms
(b) Adaptive Interface
Figure 15: Comparison of the $F_{1}$ scores on diverse OS platforms and the use of adaptive interface.
V-F2 Cross-Platform Detection Reliability: Performance Across Heterogeneous OS Environments
Figure 15(a) shows the node-level $F_{1}$ score across different OS environments. In our evaluation, E3-TRACE and E3-THEIA were conducted on a Linux system, E3-CADETS was based on a FreeBSD system, and attack scenarios #1 to #3 were evaluated on a Windows system. It can be seen that our solution benefited from the OS-specific detection mechanism, achieving superior detection performance compared to Nodlink, particularly on Linux, where system activities are highly dependent on command-line interactions. Unlike Nodlink, which employs a generic detection approach, our method adapts to system-specific behaviors by leveraging known attack signatures across different OS platforms. This tailored strategy enhances detection accuracy and reliability, ensuring more precise identification of malicious activities.
V-F3 Adaptive Detection in Dynamic Cyber Environments: Minimizing False Positives
Figure 15(b) shows the node-level $F_{1}$ score before and after the use of the adaptive interface across different OS platforms. The results of the worst-case and best-case scenarios are derived from a series of fine-tuning attempts conducted during our evaluation.
It can be observed that, AI detectors are highly sensitive to varying operational environments, necessitating dynamic fine-tuning to adapt effectively to changing conditions. Similarly, each OS environment requires fine-tuning to achieve optimal detection results.
However, most research efforts [13, 10, 11, 14, 12] fail to provide a user-friendly, adaptive interface for fine-tuning their models. Instead, they rely on pre-defined parameters within their detection mechanisms, resulting in suboptimal performance in real-world detection scenarios. Our adaptive interface effectively addresses this challenge, ensuring that our solution can be seamlessly applied in real-world detection scenarios.
(a) GPT Recommender Results on E3-TRACE
(b) GPT Recommender Results on E3-THEIA
(c) GPT Recommender Results on E3-CADETS
(d) GPT Recommender Results on Real-Time Attack Scenario #1
(e) GPT Recommender Results on Real-Time Attack Scenario #2
(f) GPT Recommender Results on Real-Time Attack Scenario #3
Figure 16: Comparison of GPT recommender results across E3 datasets and real-time attack scenarios.
(a) Progress Injection and Command-Line Evasion
(b) Living off the Land Binary (LOLBin)
Figure 17: Visualization of (a) Process Injection and Command-Line Evasion (Silent Lynx [61, 62]); (b) Living off the Land Binary (LOLBin Evilnum [65]); with GPT recommender.
(a) Advanced Command Transformation
(b) Backdoor Execution
(c) Base64 Shell Command Obfuscation
(d) Log Deletion
(e) Timestamp Forgery
(f) Firefox Extension Attack
Figure 18: Visualization of the differences in detected command-line execution chains between DefendCLI and Nodlink [13], based on the E3-series datasets and real-time evaluation.
V-F4 Enhancing the Practicability of our Solution for Actionable Attack Provenance Analysis
We replace the traditional alarm scoring framework with a GPT-powered recommender system to prioritize critical alarms, effectively reducing alarm fatigue and improving response efficiency. Figure 17 shows the visualization of some InfoPaths that are detected in our real-time evaluation with the GPT recommender. The GPT recommender first analyzes and explains the intent behind each command-line that appears potentially malicious, and then prioritizes the alarm for further examination.
Figure 16 presents the overall results of the GPT recommender across E3 datasets and real-time attack scenarios. The GPT recommender generates the top-5 malicious snapshot alarms for initial examination, while the remaining alarms are restored for further review. If the top-5 snapshot alarms exhibit malicious traces, security analysts then proceed to examine the other alarms to gather additional evidence.
TABLE X: The Scale of Attack-Related Command-Line Detected by DefendCLI and Nodlink
Dataset DefendCLI Nodlink Potential Weakness
E3-TRACE 18% 11% Obfuscation
E3-THEIA 19% 17% Context Awareness
E3-CADETS 18% 13% Low-Frequency Event
A1 15% 8% Context Awareness
A2 25% 12% Low-Frequency Event
A3 9% 3% Low-Frequency Event
V-G The Improvement of Command-Line Detection Capability
Table X presents a comparison of the scale of attack-related command-line activities detected by DefendCLI and Nodlink [13]. It can be observed that our solution demonstrates superior overall effectiveness in identifying a greater number of malicious command-line activities.
$\blacksquare$ How well does DefendCLI in the detection of obfuscation compared to Nodlink? Figures 18(a) and 18(c) illustrate the detection results of both detectors in handling advanced command-line transformations and Base64-encoded shell command obfuscation.
In Figure 18(a), the string 732F5B202F67732F5B...202F67 represents a transformation of an ASCII command into its hexadecimal form as a defense evasion technique. When converted back to ASCII, it reveals the command $s/[[:blank:]]\textbackslash{}+//g$, which is used to clean up injected spaces in obfuscated shellcode designed to evade detection. In Figure 18(c), the string 2E2F676C20262E2F67...6C2026 represents a base64 obfuscated command-line for defense evasion. When decoded the content, it reveals the command $gtcache~\&\textgreater{}/dev/null~\&$, which runs a backdoor process in the background, suppressing output and ensuring persistent session activity without logging.
Based on our evaluation results, DefendCLI successfully detects advanced command-line attacks, whereas Nodlink [13] fail to identify them. This demonstrates the superior capability of DefendCLI in handling complex command-line obfuscation techniques, ensuring more comprehensive attack detection.
$\blacksquare$ How well does DefendCLI in the attack correlation with the low-frequency event compared to Nodlink? Figures 18(b), 18(d), 18(e), and 18(f) illustrate the detection results of both detectors in handling complex attack-chain correlations involving low-frequency events.
It can be observed that DefendCLI exhibits superior overall effectiveness in identifying and correlating a broader range of relevant command-line activities, improving the analysis of the provenance of the attack.
In contrast, Nodlink [13] fails to detect the command chmod+xtcexecchmod~+x~tcexec for escalating the backdoor privileges of tcexectcexec as shown in Figure 18(b). It also does not identify clear_consoleclear\_console as the root cause of the attack as shown in Figure 18(d) and is unable to reconstruct the full sequence of attack activities due to the absence of the timestamp forgery command $date-dnull+\%s$ as shown in Figure 18(e). While it detects the crashed execution of the Firefox browser, it fails to correlate subsequent actions involving the updatemodeupdate-mode script, which signifies a deeper attack footprint related to information modification as shown in Figure 18(f).
V-H Runtime Performance
Table XI presents the execution performance of DefendCLI across diverse datasets, highlighting its efficiency through parallelization. For the E3 data sets, DefendCLI runs on a 64,GB RAM, 16-core virtual machine, processing approximately 5,000 records in 6 seconds using multiprocessing. For industrial-scale, real-time detection tasks (A1, A2, A3), the system is deployed on Microsoft Azure DataLake [66], leveraging Kafka-based streaming and ETL pipelines. Under this setup, DefendCLI achieves end-to-end response times of roughly 2 seconds, including network latency, for high-volume threat detection workloads.
VI Artifact Availability
The complete artifact package accompanying this submission is available as open-source at https://anonymous.4open.science/r/DefendCLI-6CDF. This package includes the source code for our proposed method (DefendCLI), the datasets used in our evaluations, and scripts to reproduce the experiments reported in the paper. The repository is structured to facilitate artifact evaluation; please refer to the included README file for guidance.
TABLE XI: Runtime Performance
Dataset Runtime Batch Size Platform
E3-TRACE $\approx$ 5 secs 5,000 records 64GB RAM - 16 Cores
E3-THEIA $\approx$ 5 secs 5,000 records 64GB RAM - 16 Cores
E3-CADETS $\approx$ 7 secs 5,000 records 64GB RAM - 16 Cores
A1 A2 A3 $\approx$ 2 secs / Microsoft Azure DataLake
VII Conclusion
In conclusion, DefendCLI addresses critical gaps by introducing a multilevel command-line-driven differentiation framework that enhances attack provenance analysis. By leveraging graph algorithms, explainable AI, and multi-stage correlation, DefendCLI refines provenance analysis to detect attacks existing methods often miss. Unlike traditional models relying on pre-trained normal behavior profiles, it provides contextually aware anomaly detection, improving precision while reducing irrelevant alerts.
Through an extensive evaluation of real-world APT threats, DefendCLI demonstrates effectiveness in identifying sophisticated attack chains, detecting overlooked threats, and improving actionable intelligence for security teams. By enhancing reliability, reducing detection noise, and increasing context awareness, it delivers a practical, scalable, high-precision EDR solution to strengthen cybersecurity defenses.

### v2

@@
 The results indicate that DefendCLI not only detects previously unknown attack instances, which are missed by other modern commercial solutions, but also achieves a $2.3\times$ improvement in precision over the state-of-the-art research work.
-Index Terms:
-Endpoint Detection and Response, Attack Investigation, Advanced Persistent Threat.
 I Introduction
@@
 II Related Work
-The attack provenance graph is a powerful and comprehensive tool that maps the origins, evolution, and interconnections of entities and events within a system, with a particular emphasis on data flow and processes [6, 7].
-By constructing detailed network diagrams, it traces the journey of data, illustrating how information is accessed, transmitted, and modified across various endpoints.
-We argue that, despite significant progress made by existing methods in leveraging provenance analysis for system monitoring, their effectiveness in detecting and explaining command-line activities remains limited.
-These activities, however, often provide actionable insights for threat confirmation, which are closely related to the RQs discussed above.
-This section provides a comparative analysis of related work by discussing their limitations across four key dimensions [8] that significantly impact the effectiveness of the command-line activity detection: restricted effectiveness obfuscation detection (OD), unclear and insufficient attack correlation (AC), incapable of low-frequency events detection (LFD), and limited context awareness (CA).
-Table I qualitatively compares DefendCLI with previous methods in their granularity levels to detect command-line activities, based on their respective algorithmic designs (AD).
-A more rigorous quantitative evaluation is provided in Section V for the evidential validation.
-TABLE I: Comparison of DEFENDCLI and Other Methods in Command-Line Detection (Granularity Levels). : Dominant, : Decent, : Moderate, ❍: Basic.
+This section provides a comparative analysis of related work across four key dimensions critical for command-line activity detection: obfuscation detection (OD), attack correlation (AC), low-frequency detection (LFD), and context awareness (CA) [23].
+Table I presents a feature analysis of DefendCLI relative to prior methods based on their algorithmic designs (AD).
+1
+OD: Challenges in Obfuscation Detection [6].
+Most existing approaches primarily leverage structural anomalies, a strategy that may not fully capture the semantics of obfuscated commands (e.g., polymorphic commands or encoded payloads).
+Holmes [22] provides effective detection via TTP-based rules.
+Methods like Unicorn [7], DepComm [35], ProGrapher [36], Nodlink [17], and DISTDET [5] focus largely on structural changes or statistical deviations.
+While effective for general anomalies, these systems face challenges in parsing sophisticated obfuscation techniques without semantic analysis.
+2
+AC: Granularity in Attack Correlation [37].
+Effective correlation is essential for reconstructing attack chains.
+Nodlink [17] demonstrates strong performance by employing Steiner trees for node-level correlation, while Holmes [22] utilizes TTP rules to provide a cohesive view of adversarial behavior.
+Unicorn, DepComm, and ProGrapher rely on dependency tracking, which captures structural relationships.
+DISTDET [5] emphasizes efficient log reduction and deduplication, which prioritizes scalability but may limit the resolution of complex relationships in multi-stage APTs.
+3
+LFD: Sensitivity to Low-Frequency Events [24].
+Detecting rare, non-repetitive patterns is vital for identifying novel or targeted attacks.
+Many current methods (Unicorn, DepComm, ProGrapher, Nodlink, DISTDET) utilize historical frequency profiles, which are effective for repetitive behaviors but may be less sensitive to rare events lacking precedence.
+Holmes [22] offers detection through predefined rules, whereas anomaly-based baselines in other systems must balance sensitivity with false-positive rates when encountering stealthy, low-frequency threats.
+4
+CA: Context Awareness Integration [18].
+Distinguishing benign from malicious commands often depends on execution context (e.g., user privileges, parent processes).
+DepComm [35] incorporates system attributes as dependencies to enhance context.
+Holmes [22] uses rule-based context for known attacks.
+Approaches such as Unicorn, ProGrapher, Nodlink, and DISTDET primarily focus on graph structure.
+Integrating robust environmental variables remains a key area for development to accurately differentiate between legitimate operations and malicious mimicry.
+Figure 3: Attack-Clause Sketch and Attack-Evidence Awareness: The attack-clause sketch module includes (1) provenance construction for building behavior graphs and (2) command-line evaluation to refine weights using known attacks.
+The attack-evidence awareness module handles (3) anomaly detection for unknown threats and (4) reporting via RAG-LLM.
+TABLE I: Feature Analysis of DEFENDCLI Relative to Prior Work. : Comprehensive, : Substantial, : Partial, : Limited.
 Method OD AC LFD CA AD
@@
 DefendCLI Multi-Level Deviation
-1
-OD: Restricted Effectiveness in Obfuscation Detection.
-[15] A significant limitation of many previous approaches lies in their heavy reliance on detecting structural differences or anomalies within execution graphs to identify malicious activities.
-Although this strategy can be effective for well-defined and previously observed attack patterns, it falls short when addressing the challenges posed by obfuscated command-line execution.
-Attackers often leverage obfuscation techniques to disguise malicious behavior, making it difficult for detection systems to identify these threats using traditional anomaly-based or structural methods.
-The lack of robustness against obfuscation represents a critical vulnerability in these systems, which attackers can exploit to evade detection.
-Detecting obfuscated commands is essential for mitigating advanced attack techniques that often rely on subtle modifications to otherwise legitimate processes.
-However, most existing methods struggle to achieve satisfactory results in this area.
-For example, Holmes [9] demonstrates decent effectiveness in identifying some obfuscation patterns, but its scope remains limited and fails to provide a comprehensive solution due to its rule-based nature.
-On the other hand, Unicorn [10], DepComm [11], ProGrapher [12], Nodlink [13], and DISTDET [14] show basic performance in this domain.
-These methods primarily focus on detecting more straightforward anomalies or structural changes and are not designed to handle sophisticated obfuscation techniques, such as polymorphic commands, encoded payloads, or the use of benign system utilities for malicious purposes.
-This limitation significantly reduces their effectiveness in combating modern, highly evasive attack strategies, leaving critical gaps in system security.
-2
-AC: Unclear and Insufficient Attack Correlation.
-[16] Similarly, these methods often struggle with attack correlation, as identifying relationships between commands typically requires a contextual understanding that goes beyond detecting mere structural deviations.
-Effective attack correlation is crucial for piecing together the sequence of events within an attack, enabling security teams to understand the full scope and intent of malicious activities.
-Reconstructing attack chains through command correlation is vital for analyzing complex threat patterns and attributing malicious activities to specific campaigns.
-Among existing approaches, Nodlink [13] demonstrates dominant performance by using a steiner tree framework to correlate node-level events.
-Holmes [9] also has decent success, employing attack chain reconstruction techniques that provide a cohesive view of adversarial behavior.
-Similarly, Unicorn [10], DepComm [11], and ProGrapher [12] employ graph-based or dependency-oriented methods to trace structural relationships and event dependencies.
-While effective to a degree, these methods often fall short in capturing the nuanced contextual information necessary to identify more sophisticated and stealthy attack patterns.
-In contrast, DISTDET [14] performs basically in this domain due to its lack of robust correlation mechanisms, as it simply deduplicates similar command-line entries without distinguishing between malicious and benign instances.
-Its limited ability to interpret complex relationships between commands diminishes its effectiveness in reconstructing multi-stage attack chains.
-This shortcoming is particularly problematic for detecting advanced persistent threats (APTs), which often rely on well-coordinated, multi-layered tactics to achieve their objectives.
-Overall, the inability of existing methods to fully capture and correlate command-line interactions represents a significant barrier to achieving comprehensive threat detection and analysis.
-3
-LFD: Incapable of Low-Frequency Events Detection.
-[17] Furthermore, detecting low-frequency patterns, which are often indicative of novel or highly targeted attacks, remains a critical gap in existing solutions.
-These patterns typically lack the historical context or recurrence that anomaly-based systems rely upon, making them harder to identify with conventional techniques.
-Methods that depend solely on anomaly detection tend to produce high false positive rates or miss subtle, context-dependent signals that require a deeper understanding of system behavior.
-Detecting rare or low-frequency command-line patterns is particularly vital for identifying emerging threats and sophisticated targeted attacks.
-However, most existing approaches fail to effectively address this dimension.
-For instance, Holmes [9] offers moderate rule-based support for low-frequency event detection.
-Similarly, Unicorn [10], DepComm [11], ProGrapher [12], Nodlink [13], and DISTDET [14] primarily rely on identifying common patterns based on historical profiles of normal activity.
-However, they overlook the complexities of rare or context-specific anomalies, which frequently evolve and pose significant security challenges.
-This limitation hampers their ability to detect and respond to advanced stealth attacks that exploit the blind spots of conventional detection mechanisms.
-4
-CA: Limited Context Awareness.
-[18] Finally, the lack of context awareness in many existing approaches significantly undermines their ability to differentiate between benign and malicious commands, particularly in dynamic environments where identical commands can have vastly different implications depending on their execution context.
-Context awareness is critical for accurately interpreting command behavior, as the same command might serve legitimate purposes in one situation but be part of an attack in another.
-Without the ability to incorporate and analyze contextual data, detection systems often generate high false positive rates or fail to detect subtle malicious activities embedded within otherwise benign operations.
-Context awareness enables detection systems to consider factors such as the command’s origin, associated processes, user privileges, and the broader system state, thereby reducing false alarms and enhancing detection precision.
-Among existing methods, Holmes [9] provides a moderate rule-based context awareness for known attacks, which limits its ability to distinguish between benign and unknown malicious command usage effectively.
-Similarly, Unicorn [10], ProGrapher [12], Nodlink [13], and DISTDET [14] exhibit basic performance in this area due to minimal integration of contextual data into their detection processes.
-These methods primarily rely on structural or anomaly-based techniques without considering the broader operational environment in which commands are executed.
-In contrast, DepComm [11] shows decent performance using partial contextual information within their method.
-For instance, the method may take into account some system attributes as dependencies to reduce the attack irrelevance.
-However, its reliance on context reduction often fails to capture the full spectrum of environmental variables, making them inadequate for detecting sophisticated threats that exploit contextual subtleties.
-The lack of robust context-aware mechanisms in these approaches leaves significant blind spots in detecting stealthy tactics that mimic legitimate behavior, allowing attackers to exploit contextual ambiguities and bypass traditional detection methods.
-III Detection Surface
-Figure 3 presents the detection surface of our solution, highlighting its effectiveness compared to previous research, which struggled to achieve accurate detection.
-DefendCLI focuses on three levels of detection to thoroughly monitor unusual activities, interacting with system audit logs: (1) PC: it identifies unusual system process calls across the OS-level for each endpoint; (2) CL: it analyzes each individual command-line execution within the detected system process chain; (3) NC: it provides information on network connections related to each system process.
-However, DefendCLI is not a silver bullet to detect all malicious behaviors as part of the advanced security analysis function in EDR.
-It is often necessary to combine it with other security devices, such as Anti-Virus [19] and NIDS [20], to achieve more comprehensive results.
-Our solution focuses primarily on attack activities that can be observed and monitored from system audit logs, which provide detailed behavioral and operational information.
-1
-PC: Unusual Interacted System Process Calls.
-Observing system process calls is crucial for detecting abnormal OS activities, as legitimate software typically does not invoke unrelated applications or scripts.
-Despite attackers embedding malicious code in trusted software, a technique known as Living off the Land Attack [21], process chains still provide transparency by exposing deviations from normal behavior [22].
-For instance, when malware is injected into a legitimate process, it spawns a new thread within the same namespace, altering system memory allocation.
-If a thread attempts to access another process’s memory, the system audit log records it as interprocess access, making such activities detectable through process call interactions.
-Figure 3: Detection Surface: Given two similar attack chains, previous detection methods failed to distinguish the structural differences, leading to an inability to correctly identify the mimicry attack (right) resulted in blind classification.
-2
-CL: Suspicious Command-Line Executions.
-The effectiveness of system process chains is often compromised when malware embeds itself within legitimate processes, reducing visibility and detectability.
-Our solution addresses this by analyzing command-line attributes of flagged system processes within a process chain.
-This approach, which inspects command-line inputs from both interactive shells and OS API executions, enhances security analysts’ ability to swiftly review EDR results when process chain investigations become challenging.
-Certain command-line interactions are critical for OS functions and attack phases, making them unavoidable despite evasion tactics.
-For example, the Turla APT Group [23] employs malware that leverages .NET or C++ to communicate with C2 servers, concealing malicious command-line activity.
-However, system process call analysis can still reveal anomalies, such as, Explorer.exe $\rightarrow$ Powershell $\rightarrow$ csc.exe (command-line: /noconfig…4krwc2ua.cmdline) $\rightarrow$ WMI.exe (command-line: tasklist).
-Notably, csc.exe rarely invokes 4krwc2ua.cmdline from the .NET library and should not trigger WMI.exe to execute tasklist, marking it as a potential anomaly.
-3
-NC: Rarely Seen External Network Connections.
-To boost detection capabilities, we integrate basic network data along with process-level and command-line-level analysis.
-Unlike network intrusion detection that uses detailed network attributes from packet inspection, the system audit log only records the essential network information IP: Port.
-The advantage of this method is its ability to link network connections directly with specific system processes.
-In the context of the Trula APT case, the detection outcome will be supplemented with network connections such as, Explorer.exe $\rightarrow$, Powershell $\rightarrow$, csc.exe (command-line: /noconfig/…//4krwc2ua.cmdline) $\rightarrow$, WMI.exe (command-line: tasklist) $\rightarrow$, and connection: 154.26.156[.]62:4444.
-This series reveals an unusual external connection by csc.exe, a fundamental component of .NET, to execute system commands through the 4krwc2ua.cmdline library object.
-Figure 4: Attack-Clause Sketch and Attack-Evidence Awareness: The attack-clause sketch module includes (1) provenance construction for building behavior graphs and (2) command-line evaluation to refine weights using known attacks.
-The attack-evidence awareness module handles (3) anomaly detection for unknown threats and (4) reporting via RAG-LLM.
-IV DefendCLI
-To address the above challenges, we propose DefendCLI, a comprehensive system built with significant engineering effort to support actionable threat verification.
-We begin by outlining the system architecture, then describe each component following its execution workflow.
-For each part, we explain the design rationale and discuss the techniques used to implement our solution.
-IV-A System Overview
-Figure 4 illustrates the architecture of the DefendCLI system, which adopts a highly modular design to facilitate extensibility and clarity.
-The system is composed of two core components: Attack-Clause Sketch, responsible for constructing the attack provenance graph, and Attack-Evidence Awareness, dedicated to threat detection.
-1
-Attack-Clause Sketch. The attack-clause sketch can be constructed to systematically represent and trace the logical structure of an attack as discussed below.
-IV-A1 Attack Provenance Construction
-It begins with the attack provenance construction module, which is designed to capture the interactive relationships among multi-level system events.
-This module is carried out in three key steps: (1) attack-irrelevance reduction: a reduction method is used to simplify the graph while maintaining its critical structural and relational integrity.
-(2) attack-chain causal inference: a causal inference technique is applied to identify potential attack-related edges and nodes through various weighting schemes, enhancing the graph’s interpretability and analytical value.
-(3) interphase attack association: a solution is implemented to assess potential relationships between partitioned subgraphs, ensuring that multi-staged attacks are effectively correlated.
-IV-A2 Command-Line Differentiation Evaluation Based on Known Facts
-Following the attack provenance construction phase, the system transitions to the evaluation of command-line actions.
-This phase integrates multiple evaluation strategies.
-Initially, process chains and their associated command-line executions are analyzed using Sigma rules [24].
-These rules, crafted by security experts and organizations, identify specific malicious signatures, significantly reducing false positives when malicious actions match these signatures.
-Next, undetected command-line actions are cross-referenced against a repository of defense evasion rules [25, 26], which address common system command-line patterns often exploited for evasion.
-The evaluation results are then used to adjust the weights within the attack provenance graph, improving the detection accuracy.
-2
-Attack-Evidence Awareness. The attack-evidence awareness module is detailed below, highlighting the core elements that enable effective identification and contextualization of attack-related evidence.
-IV-A3 Explainable Anomaly Detection for Unknown Attacks
-As attack clauses accumulate, the corresponding evidence becomes increasingly explicit and compelling, necessitating timely detection.
-In the explainable anomaly detection phase, the system conducts a thorough analysis of the partitioned attack provenance graph.
-This process begins with Information Path (InfoPath) traversal, which identifies and ranks anomalous paths based on weighted scoring.
-While effective for known threats, rule-based detection alone is insufficient to capture previously unseen attack behaviors.
-To address this limitation, we propose a SimHash-based similarity framework that analyzes unknown command-line instances.
-This framework informs the selection of embedding strategies that maximize intra-class separability, thereby improving contextual awareness of novel attack patterns.
-To further enhance robustness, we incorporate an unsupervised voting mechanism that aggregates outputs from multiple machine learning models, reducing individual model bias and improving detection stability.
-This integrated approach improves anomaly detection performance while maintaining interpretability.
-Additionally, adaptive hyperparameter tuning is employed to optimize system parameters in real time, ensuring consistent performance across diverse and dynamic operational contexts.
-IV-A4 Report Generation and Recommendation
-To enhance SecDevOps efficiency, we employ the RAG-GPT4o model to automate the generation of comprehensive analysis reports derived from anomaly detection outputs.
-This model systematically ranks snapshot-based recommendations, highlighting the Top-5 most suspicious activities with contextual insights.
-These actionable insights include precise interpretations of suspicious command-line activities, inferred adversarial tactics, and potential threat indicators.
-By summarizing critical anomalies concisely and contextually, security analysts can quickly evaluate whether a host is compromised, significantly reducing cognitive overload and mitigating alarm fatigue.
-This streamlined approach empowers teams to swiftly prioritize responses, enhancing overall threat management capabilities
-Figure 5: Comparison of (a) Information-Flow Focused Structure and (b) Refined Command-Line Focused Structure: Both network and command-line details are stored as memory attributes within process nodes.
-The refined structure omits file nodes to reduce redundancy and lower false alarm rates, as files are rarely the root cause [27].
-IV-B Attack-Clause Sketch
-The attack-clause module comprises two key components: (1) attack provenance construction and (2) command-line differentiation evaluation based on known attack patterns.
-These phases are critical for correlating attack clauses and quantifying the associated risk levels.
-1
-Attack Provenance Construction.
-To address the limitations of detecting malicious command-line activities, we design a refined attack provenance structure.
-This enhanced representation treats command-line inputs not merely as deviations within the global behavior graph but also as standalone semantic features.
-By incorporating command-line semantics as independent signals, the structure improves the granularity of behavioral differentiation and enhances the detection of subtle or previously unseen attacks.
-Figure 5 provides a comparative visual representation of the improvements in constructing the provenance graph.
-IV-B1 Attack Provenance Structure
-Figure 5(a) illustrates the conventional approach, termed the Information-Flow Focused Structure, as described in [28].
-This method employs a heterogeneous graph composed of diverse node types, including reverse shell processes, system processes, command-line inputs, file paths, and network activities.
-In contrast, Figure 5(b) depicts our proposed Refined Command-Line Focused Structure.
-Unlike the heterogeneous architecture, our design adopts an isomorphic graph structure with uniform node types, centering on system processes.
-Command-line and network information are encapsulated as memory attributes within each process node, reducing structural noise while preserving semantic context.
-This refined design yields multiple advantages.
-By unifying node types, it streamlines graph representation and computation, enabling more effective application of graph-based analytics such as embedding techniques, anomaly detection algorithms, and structural similarity metrics.
-The reduced complexity and consistent node attributes enhance the system’s ability to detect and differentiate subtle anomalies in attack behavior.
-IV-B2 Attack Irrelevance Reduction
-To improve the clarity and relevance of attack provenance analysis, we introduce a reduction technique that eliminates non-correlated events from the provenance graph.
-This targeted reduction is applied at both the node and edge levels, mitigating the dependency explosion problem, a well-known challenge in provenance analysis [29].
-At the node level, we remove subgraphs that do not contain any command-line elements.
-This decision is based on two key observations.
-First, normal system behavior seldom involves extensive command-line invocation.
-Second, effective incident investigation often requires the presence of command-line clauses to trace adversarial intent.
-By isolating command-line-relevant components, the proposed reduction method streamlines the graph, reduces analytical overhead, and minimizes the likelihood of misleading dependencies.
-At the edge level, a critical concern is the presence of loops and cycles within the provenance graph, which can compromise structural integrity and impede accurate data correlation.
-To address this issue, we transform the graph into a Directed Acyclic Graph (DAG) [30].
-This transformation involves systematically identifying and eliminating cyclic dependencies by removing looped edges, thereby ensuring acyclic structure and facilitating more reliable provenance analysis.
-Figure 6: Attack-chain causal inference: the process begins by calculating node weights using PageRank and Shortest-Path Betweenness Centrality.
-Subsequently, these node weights are used to determine the edge weights.
-Figure 7: Interphase attack association: a Leiden community detection method is used to correlate interphase attacks.
-IV-B3 Attack-Chain Causal Inference
-Identifying deviations that align with plausible attack inferences remains one of the most challenging problems in attack provenance analysis.
-To address this, we adopt explainable and practical techniques to isolate abnormal instances embedded within the inference process.
-Figure 6 illustrates our approach for identifying critical elements within the complete provenance graph, enabling focused and interpretable anomaly detection.
-$\blacksquare$ How can we identify the most unusual elements traversed through reachable paths?
-Attack-relevant elements are often rare and obscured by a large volume of benign activities.
-To surface these critical nodes, we leverage the PageRank algorithm [31], originally developed to rank web pages by assessing their relative importance based on link structures.
-In this context, PageRank models the provenance graph as a network, where an edge from node $A$ to node $B$ is interpreted as a “vote” indicating the significance of $B$.
-Nodes with fewer incoming connections are considered more anomalous, reflecting their rarity within the observed activity flow.
-Mathematically, the PageRank ($PR$) of a node is computed by summing the $PR$ values of all nodes linking to it, each divided by the number of outgoing links from the source node, and then scaled by a damping factor.
-This process is defined as:
-$PR(v)=(1-d)+d\sum_{u\in\text{In}(v)}\frac{PR(u)}{L(u)},$
-where $d$ is the damping factor (typically set to 0.85), $\text{In}(v)$ is the set of nodes linking to node $v$, and $L(u)$ is the number of outgoing links from node $u$.
-The algorithm operates iteratively, recalculating scores across the graph until convergence.
-Unlike its traditional application, we normalize the final $PR$ scores to reflect rareness rather than prominence.
-This inversion highlights nodes that are less connected and potentially anomalous, enabling the effective identification of suspicious or atypical attack elements within the provenance graph.
-$\blacksquare$ How can we identify the root cause of an attack among all accessible elements?
-[32] Identifying the root cause is essential for understanding attack progression.
-However, we do not define the root cause merely as the origin point in a source-to-destination path.
-Instead, we conceptualize it as a pivotal bridge node that connects early-stage activities to subsequent malicious behaviors.
-To uncover such bridge nodes, we employ the shortest-path betweenness centrality algorithm [32], which quantifies the importance of a node based on the number of shortest paths that traverse it.
-Specifically, this measure captures how frequently a node serves as an intermediary on the shortest paths between all pairs of nodes in the graph.
-In the context of attack provenance, such nodes often correspond to compromised services or applications [33] that enable lateral movement or privilege escalation.
-The computation involves the following steps: (1) Identify all node pairs $(s,t)$, where $s$ is a source and $t$ a destination.
-(2) For each pair, compute all shortest paths between $s$ and $t$.
-(3) For each path, count the occurrences of node $v$ along the path, excluding cases where $v=s$ or $v=t$.
-(4) Aggregate these counts across all pairs to obtain the betweenness centrality of node $v$.
-This metric enables the identification of structurally significant nodes that are critical to the attack’s progression, thereby supporting effective root cause analysis in complex networked environments.
-Figure 8: Explainable anomaly detection mechanism for unknown attacks: Ranked InfoPath will be used for the anomaly detection.
-SimHash recommender will examine the difference in command-line executions.
-Finally, the Police-Witness Voting Policy will select the most anomalous InfoPath for evaluation.
-Figure 9: Example of SimHashed command-line executions: the red box of $\blacksquare$ is malicious.
-$\blacksquare$ How can we integrate both causal inference methods for enhanced analysis?
-Both PageRank and shortest-path betweenness centrality evaluate the significance of nodes within the provenance graph, yielding node-level importance scores.
-To leverage the strengths of both methods within the detection framework, we transform node weights into edge weights, an essential step for enabling InfoPaths traversal during anomaly detection.
-Given a pair of nodes $m$ and $n$ connected by an edge, we compute the corresponding Edge Weight ($EW_{(m,n)}$) as the sum of their individual node scores:
-$EW_{(m,n)}=PR_{(m)}+PR_{(n)}+CB_{(m)}+CB_{(n)},$
-where $PR$ denotes the PageRank score and $CB$ represents the shortest-path betweenness centrality score of a node.
-To ensure comparability and stability across the graph, all edge weights are subsequently normalized to the range $[0,1]$.
-This normalization facilitates consistent interpretation and supports a unified, edge-centric approach to anomaly detection within the attack provenance framework.
-IV-B4 Interphase Attack Association
-Correlating multi-stage attacks within a single group poses a significant challenge, as it requires a holistic understanding of the full attack chain.
-Establishing meaningful associations between disparate phases demands both sufficient and reliable contextual information to uncover subtle dependencies across stages.
-To address this challenge, we adopt a graph partitioning strategy using community detection to group strongly connected and semantically related nodes while isolating unrelated components.
-This is commonly achieved via the Louvain algorithm [34], a modularity optimization method frequently applied in attack graph analysis [35].
-Louvain iteratively optimizes modularity until a stable partitioning is reached.
-However, a well-documented limitation of Louvain [36] is its tendency to produce arbitrarily weakly connected communities.
-In extreme cases, the resulting communities may become disjoint after successive iterations, leading to inaccurate or incomplete data extraction, as reported in Depcomm [11].
-Such fragmentation impairs the reliability of inter-phase correlation.
-To mitigate this issue, DefendCLI replaces Louvain with the Leiden algorithm [37] for community detection, as illustrated in Figure 7.
-Empirical studies [38] have shown that Leiden not only improves partition quality by preserving community cohesion but also enhances computational efficiency, enabling faster analysis of large-scale graphs.
-Furthermore, each detected community is assigned a community score, which is subsequently incorporated into the InfoPaths search process to reinforce the correlation of multi-stage attack behaviors across the provenance graph.
-2
-Command-line Differentiation Evaluation (Known Facts). Based on real-world feedback, it is inadvisable to disregard expert knowledge during early-stage detection.
-$\blacksquare$ How can we detect known attack behaviors?
-To support practical deployment, we evaluate detected system process chains using Sigma rules [24] via SIEM platforms (e.g., Elasticsearch).
-If a matching pattern is identified, the process chain is labeled as malicious and assigned a High-Risk Impact with an associated Risk Score ($RS$).
-$\blacksquare$ How can we detect known defense evasion techniques?
-Process chains are further analyzed using a curated set of evasion rules [25, 26].
-These rules adjust the weights in the attack provenance graph based on rule confidence levels.
-Evasion indicators are categorized into Medium-Risk and Low-Risk classes, each assigned a corresponding $RS$.
-$\blacksquare$ How can we integrate both methods into attack provenance analysis?
-Following command-line evaluation, identified patterns are mapped to the provenance graph using regular expressions [39].
-For each matching edge between nodes $m$ and $n$, the Refined Edge Weight ($ReEW_{(m,n)}$) is computed as:
-$ReEW_{(m,n)}=RS\times EW_{(m,n)},$
-where $EW_{(m,n)}$ is the original edge weight. Updated edges adopt the new weight, while unmatched nodes and edges retain their original values.
-IV-C Attack-Evidence Awareness
-This module comprises two primary components: (1) explainable anomaly detection to identify unknown attacks and (2) automated threat report generation with actionable insights.
-By analyzing accumulated attack clauses, these components work in tandem to enhance understanding of attack patterns and support effective threat identification.
-3
-Explainable Anomaly Detection (Unknown Attacks). To detect novel attacks based on accumulated clauses, we design an explainable anomaly detection framework, as illustrated in Figure 8.
-IV-C1 Information Retrieval (InfoPath Search)
-A central challenge in attack provenance analysis is determining how to retrieve malicious information from large volumes of provenance data using attack clauses.
-To this end, we propose InfoPath Search, a retrieval method designed to efficiently extract relevant attack-related information.
-$\blacksquare$ How can we retrieve malicious information based on attack clauses?
-InfoPaths are derived from each partitioned cluster based on interphase attack associations.
-We employ the Dijkstra algorithm [40] to compute the shortest paths between node pairs.
-Traversal begins from nodes with zero in-degree and terminates at nodes with zero out-degree, when applicable; otherwise, all paths are considered.
-This approach ensures that, in a directed provenance graph, the shortest InfoPaths reflect plausible causal attack chains from source to sink.
-$\blacksquare$ How does information retrieval relate to causal inference and command-line evaluation?
-In our framework, higher ($ReEW$) denote higher significance in causal inference.
-However, since shortest-path algorithms prioritize lower weights, we invert the edge weights using $\frac{1}{ReEW}$ to ensure that rare and anomalous paths are prioritized.
-This transformation aligns the retrieval process with the objective of uncovering unusual and potentially malicious patterns.
-$\blacksquare$ How does information retrieval support interphase attack correlation?
-Shorter InfoPaths typically indicate more suspicious activity.
-To enhance interphase attack correlation, path lengths are further adjusted using the community score (CS), favoring paths that traverse high-impact communities, that is defined as $\frac{1}{ReEW\times CS}$.
-The final InfoPath rankings are based on two factors: (i) effective path length and (ii) the diversity of system processes involved.
-This accounts for the practical observations that malware often leverages various system processes to evade detection [41].
-IV-C2 Command-Line Differentiation Evaluation Based on Unknown Facts (SimHash Method)
-To improve the identification of novel attacks, we propose a framework that delivers an impartial anomaly detection system.
-This framework bridges the gap between algorithmic uncertainty and the capability to detect previously unseen command-line activities.
-$\blacksquare$ How can we detect subtle context deviations to effectively address complex multi-stage attacks?
-Previous research often overlooks the impact of varying embedding techniques on security datasets, leading to reduced sensitivity across diverse data types.
-This oversight diminishes the granularity of detection and restricts context awareness to minor deviations, an essential limitation when addressing novel and sophisticated attacks.
-To address this issue, we utilize the SimHash [5] algorithm to randomly assess similarity within post-embedding command-line data.
-Our approach incorporates embedding techniques such as Word2Vec [42], Doc2Vec [43], and FastText [44], which have been employed in previous studies.
-This integration aims to mitigate the impact of vectorization on different types of security audit data.
-SimHash generates a unique hash value for each embedded command-line entry, as shown in Figure 9.
-The process begins by transforming command-line data into vector representations using Word2Vec [42], Doc2Vec [43], and FastText [44], which capture semantic relationships and contextual dependencies within command sequences.
-Then, a weighted hashing function is applied to generate a fixed-length binary signature for each embedded command-line entry.
-This technique ensures that similar command-line sequences produce hash values with minimal Hamming distance, whereas significantly different sequences result in more distinct hashes.
-By computing and analyzing these distances, the approach quantifies the similarity between different command-line entries.
-Finally, the embedding method that best distinguishes the differences among all evaluated command-line inputs based on SimHash similarity scores is selected for anomaly detection.
-This process ensures that vectorization effectively preserves critical distinctions within security audit data, enhancing the detection of novel attacks while reducing unnecessary false positives.
-$\blacksquare$ How can we objectively detect anomalies based on command-line behavioral differentiation?
-Recent research trends favor the use of neural networks for anomaly detection.
-However, these methods often require large volumes of high-quality training data and substantial computing resources.
-In contrast, we believe that a cost-effective approach is well-suited for attack provenance examination, which typically demands lightweight on-site deployment.
-To achieve this goal, we developed a multifactor-based detection mechanism inspired by criminal investigation methods.
-Various classifiers (IsolationForest [45], LocalOutlierFactor [46], OneClassSVM [47], EllipticEnvelope [48], KMeans [49], and GaussianMixture [50]) then respond based on their respective algorithms, ensuring diverse analytical perspectives.
-This approach mitigates algorithmic bias by preventing over-reliance on similar features.
-The final decision is determined through a binary voting system [51], where the suspect receiving the most votes is identified.
-In this context, suspects are represented by InfoPaths embedded with command-line vectors, while multiple classifiers provide supporting evidence, culminating in the voting process.
-Ultimately, InfoPaths identified as malicious and exhibiting significant command-line differentiation are selected for further analysis.
-TABLE II: Effect of Hyperparameters on Adaptive System
-Hyperparameters Positive Effect Negative Effect
-$C$ $\uparrow C\uparrow$ DR $\uparrow C\uparrow$ FAR
-RSRS $\uparrow RS\downarrow$ FAR $\uparrow RS\downarrow$ DR
-NNNN $\uparrow NN\uparrow$ DR $\uparrow NN\uparrow$ FAR
-VSVS $\uparrow VS\uparrow$ DR $\uparrow VS\uparrow$ FAR
-IV-C3 Adaptive Interface for Dynamic Environments
-Machine learning-based detection mechanisms often demonstrate limited effectiveness in real-world environments due to the dynamic nature of software and hardware ecosystems.
-Frequent updates and evolving system behavior can degrade model performance over time, necessitating retraining and continuous evaluation, both of which introduce significant human and computational overhead.
-To mitigate these challenges, we propose an adaptive interface that allows external users to adjust key hyperparameters that influence False Alarm Rate (FAR) and Detection Rate (DR), with particular emphasis on adjusting classification thresholds.
-This user-configurable approach reduces the opacity of AI-based systems and supports improved operational alignment with diverse deployment conditions.
-Through extensive experiments, we identify several hyperparameters that significantly impact system performance.
-These are summarized in Table II, and can be tuned according to local environmental characteristics.
-Automated tuning, while convenient, often leads to increased false alarms in heterogeneous settings, reinforcing the need for manual control.
-$\blacksquare$ Influence of Risk Score (RS) based on known attack facts: As described in Section IV-B4, expert knowledge is incorporated through command-line differentiation, categorizing risk into three levels: High (H), Medium (M), and Low (L), each associated with a corresponding RS.
-These default values are set as: {H = 2.5, M = 2.0, L = 1.5}, reflecting the confidence level in known attack indicators.
-$\blacksquare$ Influence of Contamination ($C$) based on audit data quality: Real-world audit data often contain inconsistencies, making it difficult to accurately distinguish between benign and malicious activity.
-To account for data imbalance, we introduce the Contamination parameter $C$, which is integrated into supported classifiers to control for class skew.
-The default value $C=0.5$ represents a moderate assumption of data imbalance.
-$\blacksquare$ Influence of Number of Neighbors (NN) and Vector Size (VS): Data intensity varies between environments, which requires flexibility in classifier configuration.
-The Number of Neighbors ($NN$) and Vector Size ($VS$) parameters help scale detection strength.
-Lower values are recommended for smaller datasets, while higher values are preferable in high-volume environments.
-The default settings are $NN=300$ and $VS=60$.
-4
-Report Generation and Recommendation.
-Alarm fatigue, a widespread problem in cybersecurity, occurs when security analysts are inundated with excessive alerts, resulting in missed critical threats and delayed responses.
-To mitigate this challenge, we introduce a snapshot alarm mechanism powered by a GPT model enhanced through a Retrieval-Augmented Generation (RAG) framework based on GPT-4o.
-The RAG system incorporates a curated set of trusted security reports [52, 53, 54], significantly improving the model’s ability to detect, contextualize, and prioritize emerging threats.
-Figure 10 illustrates the GPT model recommendation workflow.
-The mechanism ranks all suspicious InfoPaths and highlights the Top-5 for further inspection, allowing analysts to focus on the most critical alerts and enhancing situational awareness prior to full investigation.
-In addition, the model conducts an in-depth analysis of detected command-line activity, offering clear explanations of intent and potential impact.
-This contextual intelligence supports more accurate threat identification, reduces false positives, and streamlines decision-making for security operations teams.
-Figure 10: GPT-Powered Snapshot Alarm and Threat Prioritization: An overview of the GPT model’s recommendation process, which ranks suspicious InfoPaths and provides in-depth analysis of command-line activities.
-V Evaluation
+III DefendCLI
+The architecture of DefendCLI, illustrated in Figure 3, adopts a highly modular design to ensure extensibility and clarity in addressing the challenges of command-line attack detection.
+The system is fundamentally composed of two sequential computational phases: Attack-Clause Sketch, dedicated to the construction, refinement, and structural quantification of the provenance graph; and Attack-Evidence Awareness, responsible for semantic analysis, anomaly detection, and explainable threat reporting.
+III-A Attack-Clause Sketch: Graph Construction and Quantification
+This module focuses on transforming raw system telemetry into a structured, quantifiable representation suitable for causal analysis.
+It involves graph refinement, probabilistic node scoring, and interphase community association.
+Figure 4: Comparison of (a) Information-Flow Focused Structure (heterogeneous nodes) and (b) Refined Command-Line Focused Structure (isomorphic process nodes with attributes).
+The refined structure reduces graph noise by encapsulating command and network data within process nodes.
+Figure 5: Attack-chain causal inference: Node importance is calculated using PageRank (for rarity) and Shortest-Path Betweenness Centrality (for structural bridges).
+These scores are aggregated to determine initial edge weights.
+Figure 6: Interphase attack association: The Leiden community detection algorithm groups related process nodes, overcoming fragmentation issues to correlate multi-stage attacks.
+III-A1 Refined Provenance Structure and Reduction
+Traditional approaches often construct heterogeneous graphs where processes, files, sockets, and command strings are distinct nodes (Figure 4(a)).
+This results in high-dimensional, noisy graphs where semantic context is diluted.
+To address this, we define a Refined Command-Line Focused Structure (Figure 4(b)).
+Let the system execution be represented as a directed graph $G=(V,E,\mathcal{A})$.
+Here, $V$ represents the set of uniform system process nodes.
+$E\subseteq V\times V$ denotes direct causal dependencies (e.g., process forking).
+Crucially, command-line arguments and network contexts are not separate nodes but are encapsulated as an attribute set $\mathcal{A}_{v}=\{\text{cmd}_{v},\text{net}_{v},\dots\}$ resident within each process node $v\in V$.
+This isomorphic structure preserves semantic richness while significantly reducing graph complexity for downstream embedding and traversal algorithms.
+To mitigate the dependency explosion problem [10], we apply a two-stage graph reduction function $\Phi:G\to G^{\prime}$.
+First, Semantic Pruning removes nodes lacking command-line relevance.
+We define a subgraph $G_{sub}\subset G$ as irrelevant if for all nodes $v\in G_{sub}$, the command attribute $\text{cmd}_{v}=\emptyset$.
+These subgraphs are pruned as they typically represent benign background noise.
+Second, Acyclic Transformation ensures valid causal tracing by eliminating cyclic dependencies.
+$G^{\prime}$ is transformed into a Directed Acyclic Graph (DAG) via depth-first search based cycle removal, ensuring that any traversal represents a valid temporal flow of events.
+III-A2 Causal Inference via Hybrid Node Scoring
+Identifying critical nodes within massive provenance graphs requires quantifying their structural importance.
+We employ a hybrid scoring mechanism combining probabilistic ranking and centrality measures to determine initial node weights, as depicted in Figure 5.
+Quantifying Rareness (Inverted PageRank).
+We adapt the PageRank algorithm [2] to identify anomalous nodes based on topological scarcity.
+Let $PR(v)$ denote the PageRank score of node $v$.
+The iterative computation is defined as:
+$PR(v)_{t+1}=(1-d)+d\sum_{u\in\text{In}(v)}\frac{PR(u)_{t}}{L(u)}$ (1)
+where $d\approx 0.85$ is the damping factor, $\text{In}(v)$ are incoming predecessors, and $L(u)$ is the out-degree of $u$.
+Unlike standard applications where high PR indicates prominence, in attack provenance, nodes with low in-degree (lower PR) often signify rare, potentially malicious execution paths.
+We normalize these scores to emphasize rarity.
+Identifying Root Causes (Betweenness Centrality).
+To find pivotal ”bridge” nodes that connect disparate attack phases (e.g., initial access leading to privilege escalation), we utilize Shortest-Path Betweenness Centrality ($CB$) [3].
+For a node $v$, $CB(v)$ quantifies the fraction of shortest paths between all node pairs $(s,t)$ that pass through $v$:
+$CB(v)=\sum_{s\neq v\neq t}\frac{\sigma_{st}(v)}{\sigma_{st}}$ (2)
+where $\sigma_{st}$ is the total number of shortest paths from $s$ to $t$, and $\sigma_{st}(v)$ is the count of those paths passing through $v$.
+High $CB$ scores highlight structurally critical bottlenecks.
+Edge Weight Initialization.
+To facilitate path-based analysis, these node-level metrics must be projected onto the edges.
+For any directed edge $e_{mn}\in E^{\prime}$ connecting parent $m$ to child $n$, we define the initial Base Edge Weight ($W_{mn}^{base}$) as the normalized aggregate of the structural significance of both connected nodes:
+$W_{mn}^{base}=\text{Norm}\left(PR(m)+PR(n)+CB(m)+CB(n)\right)$ (3)
+where $\text{Norm}(\cdot)$ scales the resulting magnitude to the range $[0,1]$. This weight represents the inherent structural interest of the transition from $m$ to $n$.
+III-A3 Interphase Attack Association via Community Detection
+Attacks often manifest as distinct phases separated by time, making causal linkage difficult.
+To correlate these multi-stage behaviors, we partition the reduced graph $G^{\prime}$ into semantically coherent clusters.
+We utilize the Leiden algorithm [16], an improvement over Louvain that guarantees connected communities and optimizes modularity.
+As illustrated in Figure 6, Leiden partitions the graph into a set of disjoint communities $\mathcal{C}=\{C_{1},C_{2},\dots,C_{k}\}$.
+Each detected community $C_{k}$ is assigned a Community Score ($S_{C_{k}}$) based on the aggregate risk density within that cluster.
+This score serves as a coherence multiplier in subsequent analysis, reinforcing links between activities occurring within the same attack phase context.
+III-B Attack-Evidence Awareness: Detection and Reporting
+The second phase leverages the weighted provenance graph to detect threats.
+This involves differentiating command-line semantics against known signatures and employing ensemble learning for unknown anomaly detection, as outlined in Figure 7.
+Figure 7: Explainable anomaly detection mechanism: The process involves retrieving ranked InfoPaths, analyzing command-line differences via SimHash, and using a multi-classifier ”Police-Witness” voting policy for final verdict.
+III-B1 Differentiation of Known Threats (Rule-Based Boosting)
+To integrate expert knowledge, we evaluate process command-lines against curated threat intelligence, specifically Sigma rules [29] and defense evasion patterns [12].
+Let $R(e_{mn})$ be a boolean indicator function that returns 1 if the command-line context $\text{cmd}_{n}$ associated with edge $e_{mn}$ matches a known threat signature, and 0 otherwise.
+We introduce a Risk Coefficient $\alpha>1$ corresponding to the severity of the matched rule.
+The base edge weight is updated to a Refined Edge Weight ($W_{mn}^{refined}$) to reflect this deterministic knowledge:
+$W_{mn}^{refined}=\begin{cases}W_{mn}^{base}\times\alpha&\text{if }R(e_{mn})=1\\ W_{mn}^{base}&\text{otherwise}\end{cases}$ (4)
+This operation significantly amplifies the weights of paths containing known malicious indicators, ensuring their prioritization during retrieval.
+III-B2 Differentiation of Unknown Anomalies (SimHash Ensemble)
+Detecting novel attacks requires identifying subtle deviations in command-line semantics, often disguised through obfuscation.
+We propose an approach combining semantic embedding with locality-sensitive hashing.
+Semantic Hashing via SimHash.
+Standard embeddings (e.g., Word2Vec) result in high-dimensional vectors where similarity search is computationally expensive.
+To efficiently detect obfuscated variations, we apply SimHash [27].
+Let $\mathbf{v}_{c}\in\mathbb{R}^{d}$ be the dense vector representation of a command string $c$.
+The SimHash function $H:\mathbb{R}^{d}\to\{0,1\}^{f}$ maps this vector to an $f$-bit binary signature $h_{c}$.
+A crucial property of SimHash is that the Hamming distance between two signatures approximates the cosine similarity of their original vectors:
+$\text{Hamming}(h_{c_{i}},h_{c_{j}})\approx\theta(\mathbf{v}_{c_{i}},\mathbf{v}_{c_{j}})$ (5)
+This allows us to efficiently quantify semantic deviation, as shown in Figure 8, where obfuscated commands yield small but detectable Hamming distances from benign counterparts.
+Figure 8: Example of SimHashed command-line executions.
+SimHash maps semantically similar commands to nearby binary signatures, allowing detection of obfuscated variations (red box) via Hamming distance.
+Ensemble ”Police-Witness” Voting.
+To avoid algorithmic bias inherent in single-model anomaly detection, we implement an ensemble voting strategy [25].
+We employ a diverse set of unsupervised classifiers $\mathcal{M}=\{\text{iForest},\text{LOF},\text{OC-SVM},\text{EllipticEnvelope},\dots\}$.
+Each classifier $M_{j}\in\mathcal{M}$ evaluates the SimHashed InfoPaths independently.
+The final classification of a path $P$ is determined by a majority vote:
+$V(P)=\mathbb{I}\left(\sum_{M_{j}\in\mathcal{M}}\mathbb{I}(M_{j}(P)=\text{anomalous})>\frac{|\mathcal{M}|}{2}\right)$ (6)
+Paths confirmed as anomalous by the majority are flagged for analyst review.
+III-B3 InfoPath Retrieval and GPT-Powered Reporting
+The final step is to extract the most salient attack chains, termed InfoPaths, from the graph.
+An InfoPath is defined as a directed sequence of causally linked process nodes $(v_{1},v_{2},\dots,v_{k})$ representing a potential attack progression.
+To retrieve the most critical paths, we adapt Dijkstra’s shortest-path algorithm [13].
+Since standard Dijkstra minimizes cost, and our refined weights $W_{mn}^{refined}$ represent *evidence* (where higher is more significant), we define the traversal cost function as the inverse of the evidence, further weighted by the community score to favor cohesive attack phases:
+$Cost(e_{mn})=\frac{1}{W_{mn}^{refined}\times S_{C_{community(n)}}}$ (7)
+By minimizing this cost, the algorithm effectively maximizes the cumulative evidence along the path, prioritizing InfoPaths with high structural rarity, centrality, known threat indicators, and community coherence.
+To mitigate alert fatigue, the Top-$k$ retrieved InfoPaths are fed into a Retrieval-Augmented Generation (RAG) framework [26] powered by Llama-2.
+The model augments the raw path data with context from external threat intelligence reports, generating a concise, prioritized snapshot alarm that explains the attack intent and recommended actions.
+IV Evaluation
 The evaluation is divided into overall evaluation and ablation evaluation.
@@
 The ablation evaluation conducts a detailed functional analysis to assess the effectiveness of each component in our solution, providing deeper insights into its impact and performance on discussed RQs and limitations.
-V-A Attack Datasets
-V-A1 DARPA Engagement Series Datasets (Linux and Unix-Like Servers)
+IV-A Attack Datasets
+IV-A1 DARPA Engagement Series Datasets (Linux and Unix-Like Servers)
 The DARPA Transparent Computing Engagement Series (TCES) [4] are designed to support cybersecurity research by providing real-world cyber attack traces, system logs, and network activity.
@@
 Based on the attack scenarios, the E3 dataset is divided into E3-Trace, E3-THEIA and E3-CADETS.
-TABLE III: GitHub Availability, Reproducibility and Real-Time Detection of Related Work
+TABLE II: GitHub Availability, Reproducibility and Real-Time Detection of Related Work
 Method A RP RT C
@@
 DISTDET [5] Usenix Security
-TABLE IV: Licensed Commercial EDR Solutions
+TABLE III: Licensed Commercial EDR Solutions
 Tool AV WS CTI RT
@@
 Karspersky EDR [14, 15]
-TABLE V: Node-Level Results on DARPA E3 Datasets
+TABLE IV: Node-Level Results on DARPA E3 Datasets
 Method E3-Trace E3-THEIA E3-CADETS
@@
 DefendCLI 0.41 1 0.58 0.38 1 0.55 0.2 1 0.33
-TABLE VI: Attack Scenario Checklist (E3)
+TABLE V: Attack Scenario Checklist (E3)
 Dataset Attack Scenario TTPs
@@
 Figure 9: Comparison of attack-related command-line and InfoPath-Level results on E3 datasets.
-V-A2 Industrial Real-Time Detection (Microsoft Windows End-User and Servers)
+IV-A2 Industrial Real-Time Detection (Microsoft Windows End-User and Servers)
 The Industrial Real-Time Detection dataset [11] is designed to support cybersecurity research by providing real-world, real-time telemetry from virtual machines (VMs) hosted on Microsoft Azure, enabling the evaluation of real-time detection and response capabilities for defense solutions.
@@
 With realistic attack scenarios such as unauthorized access, privilege escalation, and lateral movement, this dataset is particularly valuable for advancing machine learning-driven security solutions, continuous monitoring systems, and proactive cyber defense strategies, emphasizing the importance of real-time detection and mitigation in industrial cloud environments.
-TABLE VII: Attack-TTPs and Node-Level Results on Industrial Real-Time Detection
+TABLE VI: Attack-TTPs and Node-Level Results on Industrial Real-Time Detection
 Method Attack Scenario #1 (A1) Attack Scenario #2 (A2) Attack Scenario #3 (A3)
@@
 DefendCLI (Node-Level) 0.08 1 0.15 0.38 1 0.55 0.18 1 0.31
-TABLE VIII: Attack Scenario Checklist (Industrial)
+TABLE VII: Attack Scenario Checklist (Industrial)
 Scenario APT TTPs
@@
 3 Evilnum [28] Spear Phishing Python RAT
-V-B GitHub Availability, Reproducibility and Real-Time Detection
-Table III shows the availability (A), reproducibility (RP), and real-time detection capability (RT) of related work based on big-4 security conferences (C).
+IV-B GitHub Availability, Reproducibility and Real-Time Detection
+Table II shows the availability (A), reproducibility (RP), and real-time detection capability (RT) of related work based on big-4 security conferences (C).
 The research work of DISTDET [5] and Prographer [36] are not publicly available on GitHub or other sources.
@@
 The verification of availability was conducted on 7 February 2025.
-V-C Licensed Commercial EDR Solutions
-Table IV presents the detection capabilities of commercial EDR solutions across four key security features: anti-virus (AV), web security (WS), cyber threat intelligence (CTI), and real-time (RT) protection.
+IV-C Licensed Commercial EDR Solutions
+Table III presents the detection capabilities of commercial EDR solutions across four key security features: anti-virus (AV), web security (WS), cyber threat intelligence (CTI), and real-time (RT) protection.
 For our evaluation, we obtained commercial licenses for Microsoft Defender [20], Symantec EDR [32], and Kaspersky EDR [14, 15].
 Each EDR solution was updated to the latest version, ensuring the most recent threat database was used during testing.
-V-D Attack Provenance Evaluation Criterion
+IV-D Attack Provenance Evaluation Criterion
 We use a multi-level criterion to evaluate our work, which is introduced as follows:
@@
 This allows for a direct performance comparison between our solution and commercial detection tools.
-V-E Overall Evaluation
-V-E1 Detection Results on DARPA E3 Benchmark
-Table VI presents the threat model of the three datasets in the E3 benchmark, while Table V compares node-level detection results with state-of-the-art research.
+IV-E Overall Evaluation
+IV-E1 Detection Results on DARPA E3 Benchmark
+Table V presents the threat model of the three datasets in the E3 benchmark, while Table IV compares node-level detection results with state-of-the-art research.
 Our analysis shows that while all existing methods can detect E3 attack scenarios, their node-level results vary significantly.
@@
 Figure 10: Comparison of attack-related command-line and InfoPath alarms in real-time detection.
-V-E2 Detection Results in the Industrial Real-Time Detection
-Table VIII presents the threat model for industrial real-time detection, incorporating modern APT threats relevant to 2025.
+IV-E2 Detection Results in the Industrial Real-Time Detection
+Table VII presents the threat model for industrial real-time detection, incorporating modern APT threats relevant to 2025.
 These advanced threats are designed to evade modern defense strategies, enabling a more comprehensive evaluation of our solution’s detection capabilities in complex industrial environments.
@@
 Figure 11: Comparison of attack-chain causal inference based on the number of weighted edge count versus the number of attack-related weighted edge count.
-V-F Ablation Evaluation
-V-F1 Evaluating the Interoperability of Attack Provenance Examination Through Attack-Clause Sketch and Evidence Awareness
+IV-F Ablation Evaluation
+IV-F1 Evaluating the Interoperability of Attack Provenance Examination Through Attack-Clause Sketch and Evidence Awareness
 Our solution incorporates a variety of sophisticated algorithms to enhance the detection capabilities of attack provenance analysis.
@@
 The observed disparity between the total number of weighted edges and the subset associated with attacks further validates the model’s ability to prioritize attack-relevant edges, effectively filtering out irrelevant connections and improving the interpretability of the attack provenance graph.
-TABLE IX: The Use and Not Use of Community Detection and SimHash Recommender on Node-Level $F_{1}$ Score
+TABLE VIII: The Use and Not Use of Community Detection and SimHash Recommender on Node-Level $F_{1}$ Score
 Dataset Community Detection SimHash (Blue)
@@
 However, for industrial attack scenario #3, we observed no significant difference, as all elements were reachable regardless of the use of community detection.
-Table IX shows the node-level $F_{1}$ score with or without the use of community detection.
+Table VIII shows the node-level $F_{1}$ score with or without the use of community detection.
 The experimental results can support our claim that the application of community detection improves correlation accuracy for identifying interphase attacks, leading to more precise and effective threat analysis.
@@
 However, other methods, despite yielding fewer results, failed to identify several malicious behaviors.
-Table IX presents the node-level $F_{1}$ score with and without the use of SimHash recommender.
+Table VIII presents the node-level $F_{1}$ score with and without the use of SimHash recommender.
 The results demonstrate that incorporating the SimHash recommender enhances context awareness by effectively differentiating command-line activities, leading to more precise threat analysis.
@@
 Figure 13: Comparison of the $F_{1}$ scores on diverse OS platforms and the use of adaptive interface.
-V-F2 Cross-Platform Detection Reliability: Performance Across Heterogeneous OS Environments
+IV-F2 Cross-Platform Detection Reliability: Performance Across Heterogeneous OS Environments
 Figure 13(a) shows the node-level $F_{1}$ score across different OS environments.
@@
 This tailored strategy enhances detection accuracy and reliability, ensuring more precise identification of malicious activities.
-V-F3 Adaptive Detection in Dynamic Cyber Environments: Minimizing False Positives
+IV-F3 Adaptive Detection in Dynamic Cyber Environments: Minimizing False Positives
 Figure 13(b) shows the node-level $F_{1}$ score before and after the use of the adaptive interface across different OS platforms.
@@
 Figure 16: Visualization of the differences in detected command-line execution chains between DefendCLI and Nodlink [17], based on the E3-series datasets and real-time evaluation.
-V-F4 Enhancing the Practicability of our Solution for Actionable Attack Provenance Analysis
+IV-F4 Enhancing the Practicability of our Solution for Actionable Attack Provenance Analysis
 We replace the traditional alarm scoring framework with a GPT-powered recommender system to prioritize critical alarms, effectively reducing alarm fatigue and improving response efficiency.
@@
 If the top-5 snapshot alarms exhibit malicious traces, security analysts then proceed to examine the other alarms to gather additional evidence.
-TABLE X: The Scale of Attack-Related Command-Line Detected by DefendCLI and Nodlink
+IV-G The Improvement of Command-Line Detection Capability
+Table IX presents a comparison of the scale of attack-related command-line activities detected by DefendCLI and Nodlink [17].
+It can be observed that our solution demonstrates superior overall effectiveness in identifying a greater number of malicious command-line activities.
+$\blacksquare$ How well does DefendCLI in the detection of obfuscation compared to Nodlink?
+Figures 16(a) and 16(c) illustrate the detection results of both detectors in handling advanced command-line transformations and Base64-encoded shell command obfuscation.
+In Figure 16(a), the string 732F5B202F67732F5B...202F67 represents a transformation of an ASCII command into its hexadecimal form as a defense evasion technique.
+When converted back to ASCII, it reveals the command $s/[[:blank:]]\textbackslash{}+//g$, which is used to clean up injected spaces in obfuscated shellcode designed to evade detection.
+In Figure 16(c), the string 2E2F676C20262E2F67...6C2026 represents a base64 obfuscated command-line for defense evasion.
+When decoded the content, it reveals the command $gtcache~\&\textgreater{}/dev/null~\&$, which runs a backdoor process in the background, suppressing output and ensuring persistent session activity without logging.
+Based on our evaluation results, DefendCLI successfully detects advanced command-line attacks, whereas Nodlink [17] fail to identify them.
+This demonstrates the superior capability of DefendCLI in handling complex command-line obfuscation techniques, ensuring more comprehensive attack detection.
+$\blacksquare$ How well does DefendCLI in the attack correlation with the low-frequency event compared to Nodlink?
+Figures 16(b), 16(d), 16(e), and 16(f) illustrate the detection results of both detectors in handling complex attack-chain correlations involving low-frequency events.
+It can be observed that DefendCLI exhibits superior overall effectiveness in identifying and correlating a broader range of relevant command-line activities, improving the analysis of the provenance of the attack.
+In contrast, Nodlink [17] fails to detect the command chmod+xtcexecchmod~+x~tcexec for escalating the backdoor privileges of tcexectcexec as shown in Figure 16(b).
+It also does not identify clear_consoleclear\_console as the root cause of the attack as shown in Figure 16(d) and is unable to reconstruct the full sequence of attack activities due to the absence of the timestamp forgery command $date-dnull+\%s$ as shown in Figure 16(e).
+While it detects the crashed execution of the Firefox browser, it fails to correlate subsequent actions involving the updatemodeupdate-mode script, which signifies a deeper attack footprint related to information modification as shown in Figure 16(f).
+IV-H Runtime Performance
+Table X presents the execution performance of DefendCLI across diverse datasets, highlighting its efficiency through parallelization.
+For the E3 data sets, DefendCLI runs on a 64,GB RAM, 16-core virtual machine, processing approximately 5,000 records in 6 seconds using multiprocessing.
+For industrial-scale, real-time detection tasks (A1, A2, A3), the system is deployed on Microsoft Azure DataLake [21], leveraging Kafka-based streaming and ETL pipelines.
+Under this setup, DefendCLI achieves end-to-end response times of roughly 2 seconds, including network latency, for high-volume threat detection workloads.
+TABLE IX: The Scale of Attack-Related Command-Line Detected by DefendCLI and Nodlink
 Dataset DefendCLI Nodlink Potential Weakness
@@
 A3 9% 3% Low-Frequency Event
-V-G The Improvement of Command-Line Detection Capability
-Table X presents a comparison of the scale of attack-related command-line activities detected by DefendCLI and Nodlink [13].
-It can be observed that our solution demonstrates superior overall effectiveness in identifying a greater number of malicious command-line activities.
-$\blacksquare$ How well does DefendCLI in the detection of obfuscation compared to Nodlink?
-Figures 18(a) and 18(c) illustrate the detection results of both detectors in handling advanced command-line transformations and Base64-encoded shell command obfuscation.
-In Figure 18(a), the string 732F5B202F67732F5B...202F67 represents a transformation of an ASCII command into its hexadecimal form as a defense evasion technique.
-When converted back to ASCII, it reveals the command $s/[[:blank:]]\textbackslash{}+//g$, which is used to clean up injected spaces in obfuscated shellcode designed to evade detection.
-In Figure 18(c), the string 2E2F676C20262E2F67...6C2026 represents a base64 obfuscated command-line for defense evasion.
-When decoded the content, it reveals the command $gtcache~\&\textgreater{}/dev/null~\&$, which runs a backdoor process in the background, suppressing output and ensuring persistent session activity without logging.
-Based on our evaluation results, DefendCLI successfully detects advanced command-line attacks, whereas Nodlink [13] fail to identify them.
-This demonstrates the superior capability of DefendCLI in handling complex command-line obfuscation techniques, ensuring more comprehensive attack detection.
-$\blacksquare$ How well does DefendCLI in the attack correlation with the low-frequency event compared to Nodlink?
-Figures 18(b), 18(d), 18(e), and 18(f) illustrate the detection results of both detectors in handling complex attack-chain correlations involving low-frequency events.
-It can be observed that DefendCLI exhibits superior overall effectiveness in identifying and correlating a broader range of relevant command-line activities, improving the analysis of the provenance of the attack.
-In contrast, Nodlink [13] fails to detect the command chmod+xtcexecchmod~+x~tcexec for escalating the backdoor privileges of tcexectcexec as shown in Figure 18(b).
-It also does not identify clear_consoleclear\_console as the root cause of the attack as shown in Figure 18(d) and is unable to reconstruct the full sequence of attack activities due to the absence of the timestamp forgery command $date-dnull+\%s$ as shown in Figure 18(e).
-While it detects the crashed execution of the Firefox browser, it fails to correlate subsequent actions involving the updatemodeupdate-mode script, which signifies a deeper attack footprint related to information modification as shown in Figure 18(f).
-V-H Runtime Performance
-Table XI presents the execution performance of DefendCLI across diverse datasets, highlighting its efficiency through parallelization.
-For the E3 data sets, DefendCLI runs on a 64,GB RAM, 16-core virtual machine, processing approximately 5,000 records in 6 seconds using multiprocessing.
-For industrial-scale, real-time detection tasks (A1, A2, A3), the system is deployed on Microsoft Azure DataLake [66], leveraging Kafka-based streaming and ETL pipelines.
-Under this setup, DefendCLI achieves end-to-end response times of roughly 2 seconds, including network latency, for high-volume threat detection workloads.
-VI Artifact Availability
-The complete artifact package accompanying this submission is available as open-source at https://anonymous.4open.science/r/DefendCLI-6CDF.
-This package includes the source code for our proposed method (DefendCLI), the datasets used in our evaluations, and scripts to reproduce the experiments reported in the paper.
-The repository is structured to facilitate artifact evaluation; please refer to the included README file for guidance.
-TABLE XI: Runtime Performance
+TABLE X: Runtime Performance
 Dataset Runtime Batch Size Platform
@@
 A1 A2 A3 $\approx$ 2 secs / Microsoft Azure DataLake
-VII Conclusion
+V Conclusion
 In conclusion, DefendCLI addresses critical gaps by introducing a multilevel command-line-driven differentiation framework that enhances attack provenance analysis.